### v1

# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

Abstract
We study the computation of the rate-distortion-perception function (RDPF) for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences. In this setting, the RDPF forms a convex programming problem for which we characterize the optimal parametric solutions. We employ the developed solutions in an alternating minimization scheme, namely Optimal Alternating Minimization (OAM), for which we provide convergence guarantees. Nevertheless, the OAM scheme does not lead to a direct implementation of a generalized Blahut-Arimoto (BA) type of algorithm due to the presence of implicit equations in the structure of the iteration. To overcome this difficulty, we propose two alternative minimization approaches whose applicability depends on the smoothness of the used perception metric: a Newton-based Alternating Minimization (NAM) scheme, relying on Newton’s root-finding method for the approximation of the optimal iteration solution, and a Relaxed Alternating Minimization (RAM) scheme, based on a relaxation of the OAM iterates. Both schemes are shown, via the derivation of necessary and sufficient conditions, to guarantee convergence to a globally optimal solution. We also provide sufficient conditions on the distortion and the perception constraints which guarantee that the proposed algorithms converge exponentially fast in the number of iteration steps. We corroborate our theoretical results with numerical simulations and draw connections with existing results.
I Introduction
The rate-distortion-perception (RDP) trade-off studies the relevant problem of lossy compression under perceptual constraints on the reconstructed samples, generalizing the classical rate-distortion (RD) analysis [2]. Concurrently proposed by Blau and Michaeli in [3] and Matsumoto in [4, 5], its introduction is motivated by the increasing necessity of a theoretical framework able to incorporate observations raised by a wide body of research spanning from machine learning to multimedia applications, see e.g., [6, 7, 8, 9], remarking the presence of an inherent tension between perceptual quality and fidelity of the compressed samples. In this context, perceptual quality refers to the property of a sample to appear pleasing from the human standpoint. Empirical evaluations of perceptual quality using human scoring [10, 11, 12] show that conventional fidelity measures fail to capture human preference and perception, especially in extreme compression scenarios. Hence, classical rate-distortion theory, focusing solely on the fidelity of the reconstructed samples, may not provide an adequate theoretical framework when considering “human-oriented” data sources, such as images, audio, and video.
The mathematical representation of the RDP trade-off is embodied by the rate-distortion-perception function (RDPF), which complements the classical single-letter rate-distortion function (RDF) with a divergence constraint between the source and reconstruction distributions. The additional constraint acts as a proxy of human perception, measuring the deviation from the real source distribution, also referred to as “natural scene statistic”, following similar principles behind a class of no-reference image quality metrics [13, 14]. It is worth noticing, however, that the selection of specific divergence metrics may be application-dependent and is still an active area of research.
An alternative interpretation of the divergence constraints can be found in its potential as a semantic quality metric, i.e. a quantification of the importance of the reconstructed source from the observer’s perspective [15]. For example, in [16] a comparison between the segmentation capabilities of models trained on traditionally compressed samples against compressed samples with enhanced perceptual quality through Generative Adversarial Network (GAN)-based restoration shows a remarkable improvement in segmentation performances, especially for smaller scene objects.
I-A Related Work
Since its introduction, the RDP trade-off received substantial interest from the information theory community, deriving operational characterizations in a variety of operational scenarios. Focusing on the case where infinite common randomness is available both at the encoder and the decoder, Theis and Wagner [17] provide variable-length codes for both the one-shot and asymptotic regime exploiting the properties of the strong functional representation lemma [18]. In the context of the output-constrained RDF, but also valid for the "perfect realism" RDPF case, Saldi et al. [19] provide coding theorems for the case where only finite common randomness between encoder and decoder is available. In [20], Chen et al. derives coding theorems for the asymptotic case, focusing on the differences between three operational cases defined by the availability of randomness between encoder and decoder, i.e. infinite common randomness, only private randomness, and the deterministic case.
Similar to its classical counterpart, the RDPF does not enjoy any general analytical solution. However, for specific source distributions and particular choices of the distortion and perception measures, there exist closed-form expressions, such as for binary sources subject to Hamming distortion and total variation (TV) distance [3] and Gaussian sources under mean squared-error distortion and various perception measures [21]. The available closed forms, while providing theoretical insights on the RDP trade-off, have limited applicability when considering arbitrary distortion and perception metrics. In the case of the classical RDF, the introduction of the celebrated Blahut-Arimoto (BA) algorithm [22] mitigates the absence of a general closed-form solution, allowing for the exact computation of the RDF for the discrete case. Furthermore, its versatility inspired various approximation methods for the RD in the continuous case [23, 24] and allowed it to be adapted to a variety of source coding problems, with applications from quantum information theory [25] to goal-oriented communication [26]. Also in the RDPF case, numerical solutions have been explored to some extent. Data-driven solutions have been proposed, usually employing a generative adversarial scheme minimizing a linear combination of distortion and perception metrics, see e.g., [3, 27, 28, 29]. Despite the advantage of directly optimizing the image/video codec using only source samples, these methodologies still require particular effort as they are generally highly computational and data-intensive and may suffer from a lack of generalization capabilities. Algorithmic solutions for the RDPF estimation are also present. In the case of discrete alphabets, Chen et al. in [30] cast the RDPF problem as an entropic-regularized Wasserstein barycenter problem and propose a solving method based on the Sinkhorn algorithm applicable to arbitrary distortion measures and perception measure being either a Wasserstein-type distance, the Kullback–Leibler divergence, or the TV distance. Focusing on Gaussian sources, Serra et al. in [21] designs an alternating minimization method for the computation of the Gaussian RDPF for arbitrary fidelity and perception measures, deriving analytical solutions for the so-called “perfect realism” case. Always for the “perfect realism” case, Serra et al. [31] design a solving algorithm for general multivariate continuous sources and distortion metrics, leveraging the information-geometric aspects of the constrained rate-distortion problem. To the best of our knowledge, none of the existing methodologies addresses the generic computation of the RDPF for discrete sources relying on generalizations of the classical BA algorithm.
I-B Contributions
The objective of this work is to propose a generic algorithmic approach for the computation of the RDPF, focusing on the case of discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the class of $f$-divergences.
Our results leverage the fact that the RDPF forms a convex program under mild regularity conditions on the perception constraint (i.e., convexity on the second argument), which are satisfied by the considered class of divergences. This enables us to derive a parametric characterization of the optimal solution of the RDPF (Lemma 1), which is subsequently utilized to construct an alternating minimization procedure, thereon referred to as the Optimal Alternating Minimization (OAM) scheme, for which we also establish convergence guarantees (Theorem 2). However, the resulting structure of the OAM scheme relies on a set of implicit equations in the variables of interest, thus prohibiting the direct implementation of a generic BA algorithm similar to what is already known for the classical rate-distortion theory for $\operatorname{i.i.d.}$ sources and single-letter distortions [22]. Motivated by this technical difficulty, we propose two alternative minimization approaches that solve the implementability issue and whose applicability depends on the smoothness of the considered perception function.
  • In Section III-A, we design a Newton-based Alternating Minimization (NAM) scheme where, noticing that the solution of the OAM iterate is equivalent to a root-finding problem (Lemma 3), we use Newton’s root-finding method [32] for the computation of the optimal iteration step (Theorem 3).
  • In Section III-B, we introduce a Relaxed Alternating Minimization (RAM) scheme, where we leverage a new relaxed formulation of the structure of the iterations in the OAM scheme and, subsequently, we derive necessary and sufficient conditions to ensure convergence to a globally optimal solution (Theorem 4).
In Section IV we design the algorithmic embodiment for the NAM and RAM schemes (see Algorithm 1 and Algorithm 2, respectively) and develop suitable stopping criteria for both algorithms (Theorem 5). Moreover, we also provide sufficient conditions on the structure of the distortion and the perception constraints based on which our algorithms converge exponentially fast in the number of iterations (Theorems 7 and 8). We corroborate our theoretical findings with numerical simulations (Section V), with emphasis on the case of the TV perception metric, for which we develop a smooth approximation (Lemma 8).
I-C Notation
Let $\mathbb{N}$ denote the set of natural numbers, $[a:b]\subset\mathbb{N}$ the integer interval including its extremes, $\mathbb{R}$ the set of real numbers, and $\mathbb{R}_{0}^{+}$ the set of non-negative real numbers. Given a discrete alphabet $\mathcal{X}$, we denote by $(\mathcal{X},\mathbb{B}(\mathcal{X}))$ the Borel measurable space induced by the metric, with $\mathcal{P}(\mathcal{X})$ denoting the set of probability measures defined thereon. We indicate with $\mathcal{Q}(\mathcal{X})$ the set of all transition matrices $Q$ such that $Q\cdot p\in\mathcal{P}(\mathcal{X}^{2})$ for all $p\in\mathcal{P}(\mathcal{X})$. We denote by $\mathbb{E}\left[\cdot\right]$ the expectation operator, and by $\mathbb{E}_{q}\left[\cdot\right]$ we specify the probability distribution $q$ on which the expectation operator is applied. We indicate with square brackets the functional dependency between mathematical objects, e.g. $p[h]$ and $Q[h]$ express the functional dependence of a distribution $p\in\mathcal{P}(\mathcal{X})$ or a transition matrix $Q\in\mathcal{Q}(\mathcal{X})$ on another distribution $h\in\mathcal{P}(\mathcal{X})$. We denote with $C^{n}$ the set of $n^{th}$-times differentiable functions. Given a function $f\in C^{0}$, we denote with $\partial f$ its sub-gradient [33, Definition 8.3], while, if $f\in C^{2}$, we denote by $f^{\prime\prime}(\cdot)$ the second derivative with respect to its argument. We denote by $D(\cdot||\cdot)$ a generic divergence measure whereas $D_{f}(\cdot||\cdot)$ denotes a divergence belonging to the class of $f$-divergences. Given a vector $v\in\mathbb{R}^{d}$, we indicate with $\operatorname{diag}{v}\in\mathbb{R}^{d\times d}$ the matrix with as diagonal the elements of $v$ and zeros otherwise. Given a matrix $V\in\mathbb{R}^{d\times d}$, we denote with $\operatorname{eig}(V)$ its set of eigenvalues.
II Preliminaries
We start this section by providing the formal definition of the RDPF and an overview of its operational meaning, following the works of [3, 17]. Subsequently, given their relevance in this work, we introduce the category of statistical divergences known as $f$-divergences and conclude by providing an overview of the alternating minimization methodology.
II-A Rate-Distortion-Perception Functions
We consider finite alphabet sources and stochastic encoders/decoders pairs having access to a common source of randomness and define the minimum achievable rates when we consider per-letter expected distortion and per-letter perception constraints.
We assume that we are given an $\operatorname{i.i.d.}$ sequence of $n$-length random variables ${X}^{n}\in\mathcal{X}^{n}$ that induce the probability distribution $p_{X}\in\mathcal{P}(\mathcal{X})$. Formally, a stochastic encoder $f^{n}_{E}$ is any function belonging to the set $\mathcal{F}^{n}_{E}=\{f:\mathcal{X}^{n}\times\mathbb{R}\to\mathbb{N}\}$, whereas a stochastic decoder $g^{n}_{D}$ is any function in the set $\mathcal{G}^{n}_{D}=\{g:\mathbb{N}\times\mathbb{R}\to\mathcal{X}^{n}\}$. A stochastic code is an element of $\mathcal{F}^{n}_{E}\times\mathcal{G}^{n}_{D}$. Without loss of generality, the randomness at the encoder and decoder is represented as a single real number, i.e., an infinite number of bits, and is assumed shared by the pair, i.e., common randomness.
We let $d:\mathcal{X}^{2}\to\mathbb{R}_{0}^{+}$ denote a single-letter distortion function and $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ denote a divergence measure. Moreover, we define the sets of fidelity criteria $\{\Delta_{i}\}_{i\in[1:n]}$ and $\{\Phi_{i}\}_{i\in[1:n]}$ as follows
$\Delta_{i}\triangleq\mathbb{E}_{p_{X_{i},Y_{i}}}\left[d(X_{i},Y_{i})\right],\qquad\Phi_{i}\triangleq D(p_{X_{i}}||q_{Y_{i}})$
where $\Delta_{i}$ is the expected $i^{th}$ symbol distortion and $\Phi_{i}$ is the $i^{th}$ symbol divergence with respect to the reconstructed symbol $Y_{i}$. We are now ready to introduce the definition of achievability and that of the infimum of all achievable rates.
Definition 1.
(Achievability) Given a distortion level $D\geq 0$ and a perception constraint $P\geq 0$, a rate $R$ is said to be $(D,P)$-achievable if there exists a random variable $U$ and a sequence of codes $(f_{E}^{n},g_{D}^{n})\in\mathcal{F}^{n}_{E}\times\mathcal{G}^{n}_{D}$ with
$K_{n}=f^{n}_{E}(X^{n},U),\quad Y^{n}=g_{D}^{n}(K_{n},U)$
such that, for $i=1,\ldots,n$, the joint distribution $p_{X_{i},Y_{i}}$ satisfies $\Delta_{i}\leq D$ and $\Phi_{i}\leq P$ and
$\lim_{n\to\infty}\frac{H(K_{n}|U)}{n}\leq R.$
Then, we define
[…]
Next, we give the definition of the information-theoretic characterization of the RDPF [3].
Definition 2.
(RDPF) For a given finite alphabet source $X$ with distribution $p_{X}\in\mathcal{P}(\mathcal{X})$, a single-letter distortion $d(\cdot,\cdot)$ and a divergence $D(\cdot||\cdot)$, the RDPF is characterized as follows
$R(D,P)=$ $\min_{Q_{Y|X}\in\mathcal{Q}(\mathcal{X})}I(X,Y)$ (1)
s.t. $\quad\mathbb{E}\left[d(x,y)\right]\leq D$ (2)
$\quad D(p_{X}||q_{Y})\leq P$ (3)
where $D\in[D_{\min},D_{\max}]\subseteq[0,\infty)$, $P\in[P_{\min},P_{\max}]\subseteq[0,\infty)$,
$q_{Y}=\sum_{x\in\mathcal{X}}Q_{Y|X}p_{X}\qquad\text{and}\qquad I(X,Y)=D_{KL}(p_{X}Q_{Y|X}||p_{X}q_{Y})\triangleq I(p_{X},Q_{Y|X})$
where $I(p_{X},Q_{Y|X})$ highlights the dependency on $\{p_{X},Q_{Y|X}\}$.
We stress the following technical remarks on Definition 2.
Remark 1.
(On Definition 2) Following [3], it can be shown that (1) has some useful properties, under mild regularity conditions. In particular, [3, Theorem 1] showed that $R(D,P)$ is (i) monotonically non-increasing function in both $D\in[D_{\min},D_{\max}]\subset[0,\infty)$ and $P\in[P_{\min},P_{\max}]\subset[0,\infty)$; (ii) convex if the divergence $D(\cdot||\cdot)$ is convex in its second argument.
In the sequel, we assume that in (1), the perception constraint is an $f$-divergence, i.e., $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$, which is known to be convex in both arguments [34, Lemma 4.1]. Hence in light of the discussion in Remark 1, $R(D,P)$ forms a convex programming problem.
We conclude this section by providing a theorem that connects $R_{cr}(D,P)$ with $R(D,P)$ for general alphabets.
Theorem 1.
For $D\geq 0$, $P\geq 0$, we obtain $R_{cr}(D,P)=R(D,P)$.
Proof:
See [17, Theorem 3]. ∎
II-B Statistical divergences and the family of f-divergences.
Statistical divergences are fundamental measures used in information theory and statistics to quantify the dissimilarity between probability distributions. In their general definition, a divergence on $\mathcal{P}(\mathcal{X})$ is a function $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ such that $D(p||q)\geq 0$ for all $p,q\in\mathcal{P}(\mathcal{X})$, holding with equality if and only if $p=q$. Given the scope of this work, we focus on the family of $f$-divergences, first introduced in [35] (see also [34]). This rich class of divergences is defined as follows.
Definition 3.
($f$-divergence) Let $f:(0,\infty)\to\mathbb{R}$ be a convex function with $f(1)=0$. Then the $f$-divergence $D_{f}(\cdot||\cdot)$ associated with $f$ is defined as
$D_{f}(p||q)\triangleq\sum_{x\in\mathcal{X}}q(x)f\left(\frac{p(x)}{q(x)}\right),\qquad p,q\in\mathcal{P}(\mathcal{X})$
under the assumption that
[…]
Many commonly used divergences functions belong to the class of $f$-divergences. For example,
  • KL divergence $D_{KL}(\cdot||\cdot)$, obtained by considering $f(x)=x\log(x)$,
    $D_{KL}(p||q)=\sum_{x\in\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)$
  • Jensen-Shannon divergence $D_{JS}(\cdot||\cdot)$, where $f=x\log\left(\frac{2x}{x+1}\right)+\log\left(\frac{2}{x+1}\right)$,
    $D_{JS}(p||q)=D_{KL}\left(p\Big{|}\Big{|}\frac{p+q}{2}\right)+D_{KL}\left(q\Big{|}\Big{|}\frac{p+q}{2}\right)$
  • $TV(\cdot||\cdot)$, where $f=\frac{1}{2}|x-1|$,
    $TV(P||Q)=\frac{1}{2}\sum_{x\in\mathcal{X}}|p(x)-q(x)|$
  • $\alpha$-divergence $D_{\alpha}(\cdot||\cdot)$, where $f_{\alpha}$ is parameterized by $\alpha\in\mathbb{R}$,
    $D_{\alpha}(p||q)=\sum_{x\in\mathcal{X}}q(x)f_{\alpha}\left(\frac{p(x)}{q(x)}\right)\qquad$ $f_{\alpha}(x)=\ \begin{cases}\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(\alpha-1)}\qquad\text{if }\alpha\neq 0,\alpha\neq 1\\ x\log(x)-x+1\qquad\text{if }\alpha=1\\ -\ln(x)+x-1\qquad\text{if }\alpha=0\\ \end{cases}.$
We now provide some general properties of this family of divergences. For any $f$-divergence $D_{f}(\cdot||\cdot)$, the following properties hold:
  • (Linearity) $D_{f_{1}+f_{2}}(\cdot||\cdot)=D_{f_{1}}(\cdot||\cdot)+D_{f_{2}}(\cdot||\cdot)$
  • (Joint Convexity) for any $t\in[0,1]$ and $p_{1},p_{2},q_{1},q_{2}\in\mathcal{P}$,
    $D_{f}(tp_{1}+(1-t)p_{2}||tq_{1}+(1-t)q_{2})\leq tD_{f}(p_{1}||q_{1})+(1-t)D_{f}(p_{2}||q_{2}).$
  • (Invariance) Let $\hat{f}(x)=f(x)+c(x-1)$ for $c\in\mathbb{R}$, then $D_{f}(\cdot||\cdot)=D_{\hat{f}}(\cdot||\cdot)$.
The characterization of the family of $f$-divergences here provided summarizes the properties useful for the scope of this work. For a more in-depth mathematical analysis, we refer the reader to [36].
II-C Alternating Minimization and BA-type algorithms
The alternating minimization method is a framework for the minimization of functions of two constrained variables. Consider the following optimization problem
$\min_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)$
where $\mathcal{X}$ and $\mathcal{Y}$ are two arbitrary non-empty sets and the function $f(x,y)$ satisfies $-\infty<f(x,y)\leq+\infty$ for each $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. Furthermore, we assume that, for each $x\in\mathcal{X}$, there exists $y\in\mathcal{Y}$ with $f(x,y)<+\infty$, implying that $s\coloneqq\displaystyle\inf_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)<+\infty$. Depending on the case, the existence or uniqueness of the minimizer $(x^{*},y^{*})$ such that $f(x^{*},y^{*})=s$ may also be assumed.
The goal of the alternating minimization method is to create a sequence $\{(x^{(n)},y^{(n)})\}$ such that $\lim_{n\to\infty}f(x^{(n)},y^{(n)})=s$. Under specific conditions, such sequence can be defined using the solutions of two sub-problems; for $x_{i}\in\mathcal{X}$, $h(x_{i})=\displaystyle\operatorname*{arg\,min}_{y\in\mathcal{Y}}f(x_{i},y)$ and, for $y_{i}\in\mathcal{Y}$, $g(y_{i})=\displaystyle\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x,y_{i})$. For some initial point $y^{(0)}$, we can define the $n$-th element of the sequence as:
$x^{(n)}=g(y^{(n-1)})\qquad y^{(n)}=g(x^{(n)})\qquad\text{for }n=1,2\ldots.$
Depending on the problem, various sufficient conditions for the existence and optimality of the sequence limit have been studied. For instance,
  • In [37], Csiszár and Tusnády prove that, if the sequence $\{(x^{(n)},y^{(n)})\}$ guarantees
    $f(x,y)+f(x,y^{(n-1)})\geq f(x,y^{(n)})+f(x^{(n)},y^{(n-1)})\quad\forall x\in\mathcal{X},\forall y\in\mathcal{Y}$
    referred to as Five-Point property, then the optimality of the limit is ensured.
  • In [38], Grippo and Sciandrone prove convergence of the sequence to a stationary point of $f$, under the assumption of convexity of the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ and existence of the sequence limit.
BA-type algorithms, introduced for the numerical computation of channel capacity curve [39] and RD curve [22], represent specific instances of alternating minimization algorithms [40, Chapter 9]. In fact, in their classic formulation, both problems can be expressed as constrained minimization of a convex function on the sets of marginal distributions and transition matrices, where the properties of the sets (e.g., convexity) depend on the constraints for which the problem is formulated.
III Main Results
In this section, we present the derivation of our main theoretical results. We start by providing in the following lemma the parametric characterization of the solution of the RDPF problem, obtained by casting (1) as a double minimization problem.
Lemma 1.
(Double minimization) Let $D\geq 0$, $P\geq 0$ and let $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$. Moreover, let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$ being the Lagrangian multipliers associated with constraints (2) and (3). Then (1) can be expressed as a double minimum as
$\begin{split}R(D,P)=\min_{Q_{Y|X}\in\mathcal{Q}(\mathcal{X})}\min_{h\in\mathcal{P}(\mathcal{X})}D_{KL}(p_{X}Q_{Y|X}||p_{X}h)&+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}(D_{f}(p_{X}||q_{Y})-P)\end{split}$ (4)
where $D=\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]$ and $P=D_{f}(p_{X}||q_{Y}^{*})$, with $Q_{Y|X}^{*}$ achieving the minimum.
Furthermore, for fixed $Q_{Y|X}$, the right side of (4) is minimized by
$h[Q_{Y|X}](y)=\sum_{x\in\mathcal{X}}p_{X}(x)Q_{Y|X}(y|x)$ (5)
while, for fixed $h$, the right side of (4) is minimized by
$Q_{Y|X}[h](y|x)$ $=\frac{h(y)\cdot A[q_{Y}[h]](x,y,s)}{\sum_{i\in\mathcal{X}}h(i)\cdot A[q_{Y}[h]](x,i,s)}$ (6)
where
$A[u](x,y,s)$ $=\exp\left\{-s_{D}d(x,y)-s_{P}g(p_{X}(y),u(y))\right\}$ (7)
$q_{Y}[u](y)$ $=\sum_{x\in\mathcal{X}}Q_{Y|X}[u](y|x)p_{X}(x)$ (8)
$g(x,y)$ $=f\left(\frac{x}{y}\right)-\frac{x}{y}\partial f\left(\frac{x}{y}\right).$
Proof:
See Appendix A. ∎
We remark that, although showing a close resemblance to the classical BA solution [41, Theorem 6.3.3], Lemma 1 differs from it in (6). In particular, the perception constraint (3) induces the presence of an additional exponential term, i.e., $s_{P}g(\cdot,\cdot)$. Clearly, the classical BA implicit solution can be obtained as a special case of (6) by considering $s_{P}=0$, effectively removing the perceptual constraint. The next corollary follows as a direct consequence of Lemma 1.
Corollary 1.
$R(D,P)$ in (4) can be reformulated as follows
$\begin{split}R(D_{s},P_{s})=-s_{D}D_{s}-s_{P}P_{s}+\min_{h\in\mathcal{P}(\mathcal{X})}&\quad s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)\\ &-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[q_{Y}[h]](x,y,s)\right)\end{split}$ (9)
where
$D_{s}$ $=\sum_{(x,y)\in\mathcal{X}^{2}}\frac{p_{X}(x)h^{*}(y)A[h^{*}](x,y,s)}{\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)}d(x,y)\qquad P_{s}=D_{f}(p_{X}||h^{*})$
with $h^{*}\in\mathcal{P}(\mathcal{X})$ achieving the minimum of (9).
Proof:
The proof follows by substitution of (6) in (4). ∎
The following lemma characterizes a necessary and sufficient condition to ensure that, for given Lagrangian multipliers $s=(s_{D},s_{P})$, a distribution $h^{*}\in\mathcal{P}(\mathcal{X})$ is the optimal solution of (9), i.e., $(h^{*},Q_{Y|X}[h^{*}])$ defines a point achieving the RDPF.
Lemma 2.
Let $D_{f}(\cdot||\cdot)$ be such that $f\in C^{1}(0,\infty)$ continuous and differentiable on $(0,\infty)$ and let the vector function $c[\cdot,\cdot]:\mathbb{P}(\mathcal{X})^{2}\to\mathbb{R}^{|\mathcal{X}|}$ be such that
$c[u,r](y)=\frac{}{}\sum_{x\in\mathcal{X}}\frac{p_{X}(x)A[r](x,y,s)}{\sum_{i\in\mathcal{X}}u(i)A[r](x,i,s)}.$ (10)
Then, a probability vector $h$ yields a point on the $R(D,P)$ curve via the transition matrix $Q_{Y|X}$ defined in (6) if and only if $c[h,q_{Y}[h]](y)\leq 1$ for all $y\in\mathcal{X}$, holding with equality for any $y$ for which $h(y)$ is nonzero.
Proof:
See Appendix B. ∎
Remark 2.
It can be shown that the function $c[\cdot,\cdot]$ characterizes also the relation between a distribution $h$ and the result of the functional $q_{Y}[h]$. In fact, we can verify that for all $i\in\mathcal{X}$,
$\frac{q_{Y}[h](i)}{h(i)}=\sum_{x\in\mathcal{X}}\frac{Q_{Y|X}[h](i|x)}{h(i)}p_{X}(x)=c[h,q_{Y}[h]](i).$
Using the results of Lemma 1, we now proceed to construct an alternating minimization procedure, thereon referred to as Optimal Alternating Minimization (OAM), proving its convergence to a point of $R(D,P)$.
Theorem 2.
(OAM) Let the Lagrangian multipliers $s=(s_{D},s_{P})$ with $s_{D}\geq 0$, $s_{P}\geq 0$ be given. Let $h^{(0)}$ denote any probability vector with nonzero components and let $Q_{Y|X}^{(n+1)}\equiv Q_{Y|X}[h^{(n)}]$ and $h^{(n+1)}\equiv q_{Y}[h^{(n)}]$ be functions of the current iteration $h^{(n)}$ as defined in (6) and (8), respectively. Then, as $n\xrightarrow{}\infty$, we obtain
$D(Q_{Y|X}^{(n)})\xrightarrow{}D_{s},~{}P(Q_{Y|X}^{(n)})\xrightarrow{}P_{s},~{}I(p_{X},Q_{Y|X}^{(n)})\xrightarrow{}R(D_{s},P_{s}).$
Proof:
See Appendix C. ∎
Despite being optimal, the OAM scheme does not allow the implementation of a BA-type algorithmic embodiment. The reason stands in the parametric dependencies underlying (6) and (8), as highlighted in the following remark.
Remark 3.
(Implicit Iterate) Due to the structure of the iterations in (6) and (8), an implicit dependency of $h^{(n+1)}$ on itself appears, i.e.,
$\frac{h^{(n+1)}(y)}{h^{(n)}(y)}=\sum_{x\in\mathcal{X}}\frac{p_{X}(x)\exp\left\{-s_{D}d(x,y)-s_{P}g\left(p_{X},{h^{(n+1)}},y\right)\right\}}{\sum_{i\in\mathcal{X}}h^{(n)}(i)\exp\left\{-s_{D}d(x,i)-s_{P}g\left(p_{X},{h^{(n+1)}},i\right)\right\}}=c[h^{(n)},h^{(n+1)}](y)$ (11)
showing that the updated term $h^{(n+1)}$ exists in both the left- and the right-hand side of the equation. Consequently, the structure of (11) impedes the characterization of a closed form expression of the "updated term" $h^{(n+1)}$ as a function of only the current iteration term $h^{(n)}$.
The implementation problem of the OAM scheme prompts us to find alternative ways to compute the alternating minimization iterates. We detail in the following section two different approaches to solve the OAM issue, leveraging either the numerical solution of the implicit equation or through the relaxation of the structure of the iterations.
III-A NAM scheme
The implicit definition of $h^{(n+1)}$ in (11) suggests the application of numerical methods for its approximation. To this end, we introduce a variation to the OAM scheme, referred to as NAM scheme, where $h^{(n+1)}$ is approximated at each minimization step using Newton’s root-finding method [32].
We first demonstrate that the iteration step for $h^{(n+1)}$, i.e., (11), can be cast as a root finding problem.
Lemma 3.
Let $h^{(n+1)}$ be defined as in Theorem 2 and let $T:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ be the vector function defined as
$T[h^{(n)},u](i)\triangleq u(i)-h^{(n)}(i)\cdot c[h^{(n)},u](i),~{}\forall i\in\mathcal{X}$ (12)
where $c[\cdot,\cdot]$ is defined in (10). Then, $h^{(n+1)}$ is a root of $T[h^{(n)},\cdot]$, i.e., $T[h^{(n)},h^{(n+1)}]=0$.
Proof:
The proof follows from the evaluation of (12) in $h^{(n+1)}$ and the substitution of (11) therein. ∎
The application of Newton’s method requires the existence and the invertibility of the Jacobian matrix $J_{T}$ of the functional $T$ [32, Section 10.2]. In our case, ensuring the existence of $J_{T}$ requires a more restrictive continuity assumption on the divergence measure, i.e., $D_{f}(\cdot||\cdot)$ needs to be twice differentiable in its second argument. Although this limitation reduces the generality of the NAM scheme, we note that the majority of the commonly used divergences, i.e., see Section II-B, respect this assumption. Under this restriction, the invertibility of $J_{T}$ is shown in the following lemma.
Lemma 4.
Let $T[h^{(n)},\cdot]$ be the function defined in (12) and let the divergence measure $D_{f}(\cdot||\cdot)$ be twice differentiable in its second argument. Then, the Jacobian $J_{T}:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|\times|\mathcal{X}|}$ of the functional $T[h^{(n)},\cdot]$, defined as $J_{T}[h^{(n)},u]\triangleq\left[\frac{\partial T[h^{(n)},v](i)}{\partial v(j)}\Big{|}_{u}\right]_{(i,j)\in\mathcal{X}^{2}}$, is positive definite and has the form
$J_{T}[h^{(n)},u]=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]$ (13)
where
$M[h^{(n)},u]$
$\Gamma[h^{(n)},u]$ $=s_{P}\operatorname{diag}\left[h^{(n)}(i)\cdot\frac{\partial^{2}D_{f}(p_{X}||v)}{\partial v(i)^{2}}\bigg{|}_{u}\right]_{i\in\mathcal{X}}$ (14)
$C[h^{(n)},u]$ $=\operatorname{diag}\Big{[}c[h^{(n)},u](i)\Big{]}_{i\in\mathcal{X}}.$ (15)
Proof:
See Appendix D. ∎
We are now ready to define the structure of the iteration of Newton’s root-finding method applied to the functional $T(\cdot)$, which, as shown in Lemma 3, provides an approximation of $h^{(n+1)}$.
Theorem 3.
(Newton’s method) Assume the divergence measure $D_{f}(\cdot||\cdot)$ to be twice differentiable in its second argument and let $h^{(n+1)}$ and $h^{(n)}$ be defined as in Theorem 2. Let $T[h^{(n)},\cdot]$ and $J_{T}[h^{(n)},\cdot]$ be as defined in (12) and (13), respectively. Furthermore, let the sequence $\{u^{(k)}\}_{k=1,2,\ldots}$ for some initial point $u^{(0)}\in\mathbb{R}^{|\mathcal{X}|}$ be defined as
$u^{(k+1)}\triangleq u^{(k)}-\left(J_{T}[h^{(n)},u^{k}]\right)^{-1}\cdot T[h^{(n)},u^{(k)}].$
Then, $\lim_{k\to\infty}u^{(k)}=h^{(n+1)}$.
Proof:
The proof follows by direct application of Newton’s root-finding method [32, Section 10.2], since Lemma 3 proves that the set of solutions $h^{(n+1)}$ and the set of the roots of $T(\cdot)$ coincide, while Lemma 4 proves that $T(\cdot)$ satisfies the assumption for the convergence of the method. ∎
The implementation of the NAM algorithm illustrated in Algorithm 1 is obtained by introducing the results of Theorem 3 in the OAM scheme defined in Theorem 2. However, despite solving the main technical issues of the OAM scheme, the NAM scheme imposes limitations on the choice of the perception metric. In the next section, we provide an alternative minimization scheme that circumvents these issues.
III-B RAM scheme
[…]
Theorem 4.
(RAM) Let the Lagrangian multipliers $s=(s_{D},s_{P})$ with $s_{D}\geq 0$, $s_{P}\in[0,s_{P,\max}]$ be given and define
$\hat{Q}_{Y|X}[u](y|x)$
$\hat{q}_{Y}[u](y)$ $\triangleq\sum_{x\in\mathcal{X}}\hat{Q}_{Y|X}[u](y|x)p_{X}(x)$ (16)
where $A[\cdot]$ is defined in (7) and $v[\cdot]:\mathcal{P}(\mathcal{X})\to\mathcal{P}(\mathcal{X})$ is any functional defining a probability distribution. Let $\hat{h}^{(0)}$ be any probability vector with nonzero components and let $\hat{Q}_{Y|X}^{(n+1)}\equiv\hat{Q}_{Y|X}[\hat{h}^{(n)}]$, $\hat{h}^{(n+1)}\equiv\hat{q}_{Y}[\hat{h}^{(n)}]$, and $v^{(n)}=v[h^{(n)}]$. Then, as $n\xrightarrow{}\infty$, we obtain
$D(\hat{Q}_{Y|X}^{(n)})\xrightarrow{}D_{s},~{}P(\hat{Q}_{Y|X}^{(n)})$ $\xrightarrow{}P_{s},~{}I(p_{X},\hat{Q}_{Y|X}^{(n)})\xrightarrow{}R(D_{s},P_{s})$
if $\lim_{n\rightarrow\infty}||h^{(n+1)}-v^{(n)}||=0$ with at least linear rate of convergence.
Proof:
See Appendix E. ∎
Theorem 4 enables the implementation of the alternating minimization scheme by introducing an auxiliary variable $v[h^{(n)}]$, which approximates the correct iteration $h^{(n+1)}$ while still being a function of only the current iteration of $h^{(n)}$. Nevertheless, depending on $v$, this approximation may incur restrictions on the domain of the Lagrangian multiplier $s_{P}$ that affect convergence guarantees, as we will later discuss in Section IV.
We conclude this section with the following technical remark, highlighting the differences between the NAM and RAM schemes.
Remark 4.
(NAM vs. RAM) The main advantage of NAM is the convergence for any value of the Lagrangian multipliers $(s_{D},s_{P})$, without needing any additional condition. However, the introduction of Newton’s method requires the differentiability of the perceptual metric and the introduction of additional complexity at each iteration. RAM, on the other hand, lifts the differentiability requirement and avoids the additional computational cost at each iteration, but at the expense of a potentially smaller set of $(s_{D},s_{P})$ for which the algorithm achieves the optimal solution, potentially prohibiting the computation of the complete RDP curve.
IV Algorithmic Implementation and Convergence Analysis
This section is about the algorithmic implementation of the alternating minimization schemes derived in Section III and the characterization of their convergence rate.
We start by presenting the implementation of the NAM and RAM schemes, respectively, in Algorithm 1 and 2. Subsequently, we discuss the derivation of stopping conditions suitable for both algorithms.
Algorithm 1 Newton-based Alternating Minimization (NAM)
1:source distribution $p_{X}$; Lagrangian parameters $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\geq 0$; error tolerances $\epsilon>0$, $\delta>0$; distortion measure $d(x,y)$; initial assignment $h^{(0)}$.
2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
3:while $\omega\geq\epsilon$ do
4:     $h^{(n+1)}\leftarrow$ Newton Approximation( $p_{X}$, $h^{(n)}$, $s$, $\delta$)
5:     $c^{(n)}\leftarrow c\left[h^{(n)},h^{(n+1)}\right]$
6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}h^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
7:     $n\leftarrow n+1$
8:end while
[…]
Algorithm 2 Relaxed Alternating Minimization (RAM)
1:source distribution $p_{X}$; Lagrangian multipliers $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\in[0,s_{P,\max}]$; error tolerance $\epsilon>0$; divergence measure $D_{f}(\cdot||\cdot)$; distortion measure $d(\cdot,\cdot)$; initial assignment $\hat{h}^{(0)}$.
2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
3:while $\omega\geq\epsilon$ do
4:     $c^{(n)}\leftarrow c[\hat{h}^{(n)},v^{(n)}]$
5:     $\hat{h}^{(n+1)}\leftarrow\hat{h}^{(n)}\cdot c^{(n)}$
6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}\hat{h}^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
7:     $n\leftarrow n+1$
8:end while
[…]
Stopping Criterion
We first focus on the derivation of a stopping criterion for the RAM case, as the NAM case can be obtained by fixing the auxiliary variable $v[h^{(n)}]=q_{Y}[h^{(n)}]$ in Theorem 4, i.e., recovering the original OAM iterates. For this purpose, we need the following theorem in which we derive bounds on the RDPF.
Theorem 5.
(Bounds on RDPF) Let $\hat{Q}_{Y|X}$ and $\hat{h}$ be defined as in Theorem 4, $c()$ be as defined as in Lemma 2, and $c_{\max}=\max_{y\in\mathcal{X}}c[u,\hat{q}_{Y}[u]](y)$. Then, at the point $D=\mathbb{E}_{\hat{Q}_{Y|X}}\left[d(x,y)\right]$, and $P=D_{f}(p_{X}||\hat{q}_{Y})$, the following bounds hold
$R(D,P)$ $\geq R_{L}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]-\log(c_{\max})$ (17)
$\begin{split}R(D,P)&\leq R_{U}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\sum_{y\in\mathcal{X}}u(y)c[u,\hat{q}_{Y}[u]](y)\log(c[u,\hat{q}_{Y}[u]](y))\end{split}$ (18)
where $\hat{W}{}[\hat{h}]$ is given by
$\begin{split}\hat{W}[u]&=-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}u(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{u(y)}\partial f\left(\frac{p_{X}(y)}{u(y)}\right)\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u]\left(f\left(\frac{p_{X}(y)}{q_{Y}[u](y)}\right)-f\left(\frac{p_{X}(y)}{u(y)}\right)\right)\Bigg{]}.\end{split}$ (19)
Proof:
See Appendix F. ∎
Leveraging the bounds in (17) and (18), we can estimate the precision of the estimation of $R(D,P)$ at the $n$-th iteration by considering the estimation error $\omega=R_{U}[\hat{h}^{(n)}](D,P)-R_{L}[\hat{h}^{(n)}](D,P)$, as implemented in line 5 of both Algorithm 1 and 2.
IV-A Asymptotic Convergence Rate Analysis
In this section, we characterize the asymptotic convergence rate of the proposed minimization schemes. We start with the analysis of the convergence rate of the OAM scheme, which, although not directly implementable, serves as a reference for the characterization of the convergence rate of both the NAM and RAM schemes.
OAM Convergence Rate
We remark that the iteration structure in Theorem 2, i.e., $h^{(n+1)}$ = $q_{Y}[h^{(n)}]$ given the current iteration $n$, can be described as an implicit vector function $S:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ with $S[h](i)=h(i)\cdot c[h,S[h]](i)$, such that $h^{(n+1)}=S[h^{(n)}]$. The results of Lemma 2 characterize a distribution $h^{*}$ achieving the RDPF as a fixed point of $S(h)$, i.e. $h^{*}=S[h^{*}]$, since $c[h^{*},S[h^{*}]](i)=1,~{}i=1\ldots,|\mathcal{X}|$. Under these observations, we can analyze the convergence rate of the OAM scheme following similar steps as in [42].
The first order Taylor expansion of $S[h]$ around a fixed point $h^{*}$ is defined as
$S[h]=S[h^{*}]+J[h^{*}]\cdot(h-h^{*})+o(||h-h^{*}||)$
where $J[h]$ is the Jacobian matrix of $S[h]$ with entries $J[h](i,j)\triangleq\frac{\partial S[h](i)}{\partial h(j)},(i,j)\in\mathcal{X}^{2}$. The next theorem provides the functional form of the Jacobian for the case of Theorem 2.
Theorem 6.
(Jacobian form) The Jacobian $J(h)$ computed at the fixed point $h^{*}$ is given as
$J[h^{*}]=\left(I-M^{*}\right)\left(I-\Gamma^{*}J[h^{*}]\right)$ (20)
where $M^{*}=M[h^{*},h^{*}]$ and $\Gamma^{*}=\Gamma[h^{*},h^{*}]$ as defined in (LABEL:matrix:_Newton_M) and (14), respectively.
Proof:
See Appendix G. ∎
Next, we introduce two lemmas, in which we use the structure of (20) to identify properties of matrix $M^{*}$.
Lemma 5.
Let $\{\lambda_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the set of eigenvalues of $M^{*}=M[h^{*},h^{*}]$. Given a distortion function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ that induces a full-rank matrix $D=[e^{-s_{D}d(i,j)}]_{(i,j)\in\mathcal{X}^{2}}$, then $\lambda_{i}>0$,$\forall i\in[1:|\mathcal{X}|]$, i.e., $M^{*}$ has only positive eigenvalues.
Proof:
See Appendix H. ∎
Remark 5.
(On Lemma 5) We note that a popular example that satisfies the assumptions imposed on Lemma 5 is the Hamming distortion denoted hereinafter as $d_{H}$ [43].
Lemma 6.
Let $\{\lambda_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the set of eigenvalues of $M^{*}=M[h^{*},h^{*}]$. Then, we have that $\lambda_{i}\leq 1,\forall i\in[1:|\mathcal{X}|].$
Proof:
See Appendix I. ∎
Using Lemmas 5 and 6, we can characterize the interval that contains the set of eigenvalues of $J[h^{*}]$ and subsequently the convergence rate of Theorem 2.
Theorem 7.
(Convergence rate of Theorem 2) Let $\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the eigenvalues of $J[h^{*}]$. Then,
$1\displaystyle 0\leq\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}<$ 1 .
Moreover, let $\gamma\in[\theta_{\max},1)$. Then, there exists $\delta>0$ and $K>0$ such that if $h^{(0)}\in\{h:||h-h^{*}||\leq\delta\}$, we obtain
$||h^{(n)}-h^{*}||<K\cdot||h^{(0)}-h^{*}||\cdot\gamma^{n}$ (21)
i.e., the iterations converge exponentially.
Proof:
See Appendix J. ∎
Summarizing, under the structural constraints on the distortion function $d$ reported in Lemma 5, the exponential convergence of the OAM scheme is guaranteed by Theorem 7.
NAM Convergence Rate
The convergence rate of the NAM scheme is immediate from the OAM scheme analysis, given the close relation between the two schemes. Since the only difference lies in the introduction of Newton’s root-finding method for the estimation of the optimal iteration step, the NAM scheme enjoys the same convergence rate in terms of the number of iterations as the OAM scheme, i.e., an exponential convergence $\mathcal{O}(e^{-n})$ under the assumptions of Lemma 5. However, the added complexity from the application of Newton’s method at each iteration increases the overall iteration complexity, due to the at least linear convergence rate $\mathcal{O}(\frac{1}{m})$ of the root approximant. Therefore, the total complexity is approximately $\mathcal{O}(\frac{e^{-n}}{m})$, where $n$ and $m$ depend on the error tolerances $\epsilon$ and $\delta$ given as input in Algo. 1.
RAM Convergence Rate
Following similar steps that led to Theorem 7, the Jacobian $\hat{J}(h^{*})$ associated with the iteration scheme in Theorem 4 is characterized as
$\hat{J}[h^{*}]=(I-M[h^{*},h^{*}])(I-\Gamma[h^{*},h^{*}])$
where $M$ and $\Gamma$ are given by (LABEL:matrix:_Newton_M) and (14), respectively. Unlike Theorem 2, where the structure of (20) bounds its own eigenvalues, in this case, we need to bound the Lagrangian multiplier $s_{P}$, hence matrix $\Gamma$, to guarantee exponential convergence of the algorithm. This is proved in the following theorem.
Theorem 8.
For a given $s_{D}\geq{0}$, let $I_{s_{P}}=[0,s_{P,\max}]$ be the domain of $s_{P}$, $\{\theta_{a,i}\}_{i\in\mathcal{X}}$ the set of eigenvalues of $\hat{J}(h^{*})$ and $\theta_{\max}$ the maximum eigenvalue of $J(h^{*})$ in (20). Define the set $I_{s_{P}}^{\epsilon}=[0,s_{P,\max}-\epsilon]$ for $0<\epsilon<s_{P,\max}$. Then, there exists an $\epsilon^{\prime}$ such that if $s_{P}\in I_{s_{P}}^{\epsilon^{\prime}}$ then $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1.
Proof:
See Appendix K. ∎
Theorem 8 guarantees exponential convergence for Theorem 4 only for $s_{P}\in I_{s_{P}}^{\epsilon}$ which means that we can consider $P\in[P_{\min}(\epsilon),P_{\max}]$, depending on the characteristics of the input in a specific problem.
V Numerical Results
In this section, we validate our theoretical findings with simulations. In particular, we consider the computation of the RDPF under the NAM and RAM schemes, respectively using Algorithms 1 and 2.
V-A RDPF Computation - NAM scheme
Example 1.
Suppose that ${\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(\cdot||\cdot)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{4}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$. In Fig. 1, we display the $R(D,P)$ estimates resulting from Algorithm 1 for each divergence metric.
(a) $D_{JS}(\cdot||\cdot)$
(b) $D_{KL}(\cdot||\cdot)$
(c) $D_{\chi^{2}}(\cdot||\cdot)$
(d) $D_{\alpha=-1}$
(e) $D_{\alpha=\frac{1}{4}}$
(f) $D_{\alpha=\frac{1}{2}}$
Figure 1: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
We observe that all the computed RDPFs share similarities in the structure of the operating regions on the $(D,P)$ plane. Taking as a reference the plot of Fig. 1(a), we distinguish three cases:
  • Case I, where the perception level $P$ is not met with equality. In this region, the RDPF is equal to the RDF.
  • Case II, where, due to the distortion level $D$ not met with equality, the RDPF function is identically zero.
  • Case III, where both the distortion level $D$ and perception level $P$ are met with equality.
The results show that Algorithm 1 can completely cover the Case III region. The boundary between Cases I-III (black line) representing the RDF is obtained by setting the Lagrangian multiplies $s_{P}=0$. Furthermore, the Case I region can be obtained by extension of the relative boundaries, since all the $(D,P)$ points share the solution given by the RD problem.
V-B RDPF Computation - RAM scheme
Example 2.
Suppose that ${\cal X}=\hat{\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(p_{X}||h)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=TV(\cdot||\cdot)$. In Fig. 2, we display the $R(D,P)$ estimates resulting from Algorithm 2 for each divergence metric.
(a) $D_{JS}(\cdot||\cdot)$
(b) $D_{KL}(\cdot||\cdot)$
(c) $D_{\chi^{2}}(\cdot||\cdot)$
(d) $D_{\alpha=-1}(\cdot||\cdot)$
(e) $D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$
(f) $TV(\cdot||\cdot)$
Figure 2: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
Remarks similar to those in Example 1 can be made regarding the operating regions of the computed RDPFs. In addition, in Fig. 2(f) we compare the theoretical results of [3, Equation 6] with the numerical results obtained using Algorithm 2. This case is of particular interest due to the non-differentiability of the $TV$ distance. We observe that Algorithm 2 achieves exactly the theoretical solution of [3, Equation 6] as long as $D\leq{D}_{\max}=0.15$. We attribute this behavior to the limitations in the domain of the Lagrangian multiplier $s_{P}$, as discussed in Theorem 8. We address this issue in detail in the following section.
V-C On the convergence under TV perception constraint
As observed in Example 2, in the case of TV distance the RAM algorithm is not able to converge to proper solutions in the region $\Omega=\{(D,P)\in\mathbb{R}_{+}^{2}:D\geq p=0.15\}$. We argue that the problem stems from the values of the Lagrangian multipliers $s=(s_{D},s_{P})$ associated with $\Omega$ requires values of $s_{P}$ that do not guarantee the convergence of the algorithm, as reported in Theorem 8.
To solve the issue, we propose an approximation of the $TV(\cdot||\cdot)$ distance through a sequence of $f$-divergences $\{D_{f_{n}}\}_{n\in\mathbb{N}}$ such that ${D_{f_{n}}}\to TV$ for $n\to\infty$. We start with the following general property.
Lemma 7.
For a divergence metric $D_{f}(\cdot||\cdot)$, let the set $\mathcal{L}_{D_{f}}(D,P)$ be defined as
$\mathcal{L}_{D_{f}}(D,P)\triangleq\{Q_{Y|X}\in\mathcal{Q}:\mathbb{E}_{Q_{Y|X}}\left[d(x,y)\right]\leq D\land D_{f}(p_{X}||q_{Y})\leq P\}.$
Given $D_{f}$, $D_{g}$ divergence metrics with $D_{f}(p||q)\leq D_{g}(p||q),\forall p,q\in\mathcal{P}$, then $\mathcal{L}_{D_{g}}(D,P)\subseteq\mathcal{L}_{D_{f}}(D,P)$. Moreover, for the associated RDPF problems
$R_{D_{f}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{f}}(D,P)}I(p_{X},Q_{Y|X})\quad\text{and}\quad R_{D_{g}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{g}}(D,P)}I(p_{X},Q_{Y|X})$
the inequality $R_{D_{g}}(D,P)\geq R_{D_{f}}(D,P)$ holds.
Proof:
The inequality $R_{D_{g}}(D,P)\geq R_{D_{f}}(D,P)$ holds if $\mathcal{L}_{D_{g}}\subseteq\mathcal{L}_{D_{f}}$, which is a trivial implication of $D_{f}(p||q)\leq D_{g}(p||q),\forall p,q\in\mathcal{P}$. ∎
We can now characterize a sequence $\{D_{f_{n}}\}$ such that ${D_{f_{n}}}\to TV$ for $n\to\infty$ and ${D_{f_{n}}}\leq TV$.
Lemma 8.
Let ${f_{n}}$ be the sequence of convex functions defined as $f_{n}(x)=\frac{2}{\pi}(x-1)\arctan(n(x-1))$ and let $\{D_{f_{n}}(\cdot||\cdot)\}_{n=1,2,\ldots}$ be the sequence of associated $f$-divergences. Then, for $n\to\infty$, $D_{f_{n}}\to TV$ uniformly. Furthermore, for $n=1,2,\ldots$ and for all $p,q\in\mathcal{P}(\mathcal{X})$, $D_{f_{n}}(p||q)\leq TV(p||q)$.
Proof:
See Appendix L. ∎
The results from Lemmas 7 and 8 guarantee that, $\forall n\in\mathbb{N}$, the RDP problem defined for the perception metric $D_{f_{n}}(\cdot||\cdot)$ act as lower-bound for the RDP problem defined with $TV(\cdot||\cdot)$ perception metric. Furthermore, since $D_{f_{n}}(\cdot||\cdot)$ is a smooth function, we can apply NAM scheme to compute the associate RDPF. Fig. 3 shows the RDPFs for a source $X\sim Ber(0.15)$ under Hamming distortion and perception measure $D_{f_{n}}$, for $n=\{1,10,100\}$. As expected, for increasing $n$, the $D_{f_{n}}$ perception metric becomes a better approximation of the $TV(\cdot||\cdot)$ perception.
(a) $TV(\cdot||\cdot)$
(b) n = 1
(c) n = 10
(d) n = 100
Figure 3: Comparison between the RDPF under $TV$ perception, computed with RAM scheme, and the RDPF under $D_{f_{n}}$, computed with NAM scheme, for $n$={1,10, 100}.
VI Conclusion and Future Directions
In this paper, we studied the computation of the rate-distortion-perception function for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences. We showed that the introduction of the perception constraint induces a non-trivial generalization of the classical optimality conditions in terms of reconstruction distribution and transition matrix $(h^{*},Q_{Y|X}^{*})$, resulting in the OAM scheme. The presence of implicit equations in the formulation of the OAM scheme, implying the impossibility of direct algorithmic implementation, led to the formulation of alternative minimization schemes, NAM and RAM, each with its applicability conditions. Furthermore, we analyzed the asymptotic complexity of the proposed schemes and provided numerical results to validate our theoretical findings.
As an ongoing research, we plan to extend the derived optimality conditions to general measure spaces, as well as to study the behavior of the proposed algorithm when applied to continuous sources.
Appendix A Proof of Lemma 1
The double minimization follows immediately from properties of the KL divergence, reported in [41, Theorem 5.2.6], which allows to rewrite (1) as (4). Moreover, the characterization (5) of the minimizer $h$ follows from the same properties.
On the other hand, the minimizing $Q_{Y|X}$ can be characterized by applying the Karush-Kuhn-Tucker (KKT) conditions [44] on the dual formulation of (1). The Lagrangian associated with the primal problem has form:
$\begin{split}L(Q_{Y|X},s,\lambda,\mu)&=D_{KL}\left(p_{X}Q_{Y|X}||p_{X}h\right)+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}\left(D_{f}(p_{X}||q_{Y})-P\right)\\ &\quad+\sum_{x\in\mathcal{X}}\lambda_{x}\left(1-\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)\right)-\sum_{(x,y)\in\mathcal{X}^{2}}\mu_{x,y}Q_{Y|X}(y|x)\end{split}$
where the last two sets of constraints refer to the positivity and normalization constraints on $Q_{Y|X}$. Differentiating $L(Q_{Y|X},s,\lambda,\mu)$ with respect to the primal variables $Q_{Y|X}(y|x),\forall(x,y)\in\mathcal{X}^{2}$, we obtain:
$\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}+s_{D}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}+s_{P}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}+\lambda_{x}-\mu_{x,y}\end{split}$
where:
$\begin{split}\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}&=p_{X}(x)\left(\log(\frac{Q_{Y|X}(y|x)}{h(y)})+1\right)\end{split}$
$\begin{split}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}&=p_{X}(x)d(x,y)\end{split}$
$\begin{split}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}&=p_{X}(x)g(p_{X},q_{Y},y).\end{split}$
Enforcing stationarity and complementary slackness
$\begin{cases}\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=0\\ \mu(x,y)Q_{Y|X}(y|x)=0\\ \end{split}\end{cases}$ (22)
we solve for $Q_{Y|X}(y|x)$ while choosing $\lambda(x)$ such that $\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)=1$, $\forall x\in\mathcal{X}$, resulting in (6) and (7). This completes the proof.
Appendix B Proof of Lemma2
Using the results of Corollary 1, we can apply KKT conditions to (9). Thus, a minimum for $h$ must satisfy:
$\begin{split}\frac{\partial}{h(y)}\Bigg{[}-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[h](x,y,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)+\lambda\sum_{y\in\mathcal{X}}h(y)\Bigg{]}\geq 0\end{split}$
which reduces to
$\lambda-c[h,q_{Y}[h]](y)\geq 0.$
The Lagrangian multiplier $\lambda$ is evaluated by multiplying by $h$ and summing over $y\in\mathcal{X}$, giving $\lambda=1$. This concludes the proof.
Appendix C Proof of Theorem 2
Let $\textbf{h}=(h^{(0)},h^{(1)},\ldots)$, and $\textbf{Q}=(Q_{Y|X}^{(1)},Q_{Y|X}^{(2)},\ldots)$ be the sequences of probability vectors and transition matrices obtained by the chain of alternating minimization $h^{(n)}\xrightarrow{}Q_{Y|X}^{(n+1)}\xrightarrow{}h^{(n+1)}$. Let $A^{(n)}=A[h^{(n)}]$ and define the functionals $V[h^{(n)}]$ and $W[h^{(n)}]$ as
$\begin{split}&V[h^{(n)}]=D_{KL}\left(p_{X}\cdot Q_{Y|X}^{(n+1)}||p_{X}\cdot h^{(n)}\right)+s_{D}\mathbf{E}_{Q_{Y|X}^{(n+1)}p_{X}}[d(x,y)]+s_{P}D_{f}(p_{X}||h^{(n+1)})\end{split}$ (23)
$\begin{split}&W[h^{(n)}]=s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h^{(n)}(y)A^{(n)}(x,,s)\right).\end{split}$ (24)
Using Theorem 1 and Corollary 1, we can show that $V[\textbf{q}]$ is non-increasing on the sequence q by first fixing $h^{(n)}$ and minimizing over Q (resulting in $Q_{Y|X}^{(n+1)}$) and then fixing $Q_{Y|X}^{(n+1)}$ and minimizing over q (resulting in $h^{(n+1)}$). $W[h^{(n+1)}]$ is the result of the minimization over Q between $V[h^{(n)}]$ and $V[h^{(n+1)}]$, resulting in a non-increasing sequence $V[h^{(n)}]\geq W[h^{(n)}]\geq V[h^{(n+1)}]$. Given $V[\textbf{q}]$ non-increasing and bounded from below, $V[\textbf{q}]$ converges to some number $V^{\infty}$.
Let now $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and let
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)\\ &\quad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{\left(n\right)}(x,i,s)\right)\end{split}$ (25)
where
$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)$ $=-s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]-s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
$\quad+s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right).$
We can introduce an upper bound to (25) by noticing that
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}&p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}\cdot h^{*}}{h^{(n)}\cdot Q_{Y|X}^{*}}\right)\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{n}\log\left(\frac{h^{*}}{h^{(n)}}\right)-1=0.\end{split}$ (26)
As a result, we obtain
$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)\leq$
$\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)+s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]+s_{P}D_{f}(p_{X}||h^{*})$
$\quad+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)\right]$
$\quad-\Biggl{[}-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{(n)}(x,i,s)\right)+s_{P}\sum_{y\in\mathcal{X}}h^{n+1}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}$
$\leq V[h^{*}]-W[h^{(n)}]+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}(p_{X}||h^{*})\right]$
$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}.$
Notice that for any iteration $n\geq 0$, we have $W[h^{(n)}]\geq V[h^{(n+1)}]\geq V[h^{*}]$ and, subsequently, the following inequalities
$0\leq W[h^{(n)}]-V[h^{*}]\leq G[h^{(n)}]$ (27)
where
$G[h^{(n)}]$ $=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)+s_{P}\left\{\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)\right\}$
$\quad+s_{P}\Biggl{\{}\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]\Biggr{\}}.$
Since $s_{P}\geq 0$ and due to the fact that
$\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)+\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]$
$=D_{f}(p_{X}||h^{(n+1)})-D_{f}(p_{X}||h^{(*)})-\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
$\quad+\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)\frac{p_{X}(i)}{h^{(n+1)}}\partial f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
$=D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||h^{(*)}\right)-\sum_{i\in\mathcal{X}}\left(h^{(n+1}(i)-h^{*}(i)\right)\frac{\partial D_{f}(p_{X}||u)}{\partial u(i)}\Bigg{|}_{h^{(n+1)}}$
$=-\left(D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*})\right)\stackrel{{\scriptstyle(a)}}{{\leq}}0$
where $T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*}))$ is the first order Taylor expansion of $D_{f}(p_{X}||\cdot)$, centered in $h^{(n+1)}$ and evaluated in $h^{(*)}$, and $(a)$ is verified since $D_{f}(p_{X}||\cdot)$ is a convex function in its second argument. Therefore $\forall n\in\mathbf{N}$, we obtain
$W[h^{(n)}]-V[h^{*}]\leq$ $\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right).$
Summing over $N$ terms, we obtain
$\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ $\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)$
$=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(N+1)}}{Q_{Y|X}^{(1)}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{Q_{Y|X}^{(1)}}\right)$
where $(b)$ follows using the logarithm inequality Since at any iteration $n$, $W[h^{(n)}]-V[h^{*}]\geq 0$ and for all integers $N>0$ the partial sum is upper-bounded by a constant $L(h^{*},h^{(0)})$ dependent only on the initial probability assignment $h^{(0)}$, we have that $\lim_{N\to\infty}\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ exists and it is finite hence $\lim_{n\to\infty}W[h^{(n)}]-V[h^{*}]=0$. This completes the proof.
Appendix D Proof of Lemma 4
We first derive the functional form of $J_{T}$:
By defining matrices $M$, $\Gamma$ and $C$, respectively, as in (LABEL:matrix:_Newton_M), (14), (15), the matrix $J_{T}(\cdot)$ can be rewritten as in (13). To prove its invertibility, we need to ensure that 00 is not part of the set of eigenvalues of $J_{T}(y)$, i.e., $0\notin\operatorname{eig}(J_{T}(y)),~{}\forall y\in\mathbb{R}^{|\mathcal{X}|}$. Noticing that $M_{i,j}\geq 0,\forall(i,j)\in\mathcal{X}^{2}$ and $\sum_{i\in\mathcal{X}}M_{i,i}(y)=C_{i}(y)$, we can define the sets $D_{i}$ as:
$D_{i}=\{\lambda\in\mathbb{R}:\left|\lambda-\left(C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\right)\right|\leq C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\}$ (28)
Since $\bigcup_{i\in\mathcal{X}}D_{i}\subseteq\mathbb{R}_{0}^{+}$, we can apply Gershgorin Circle Theorem [45] to prove that $C[h^{(n)},u]-M[h^{(n)},u]$ has only non-negative eigenvalues. Moreover, since $D(p_{X}||\cdot)$ is a convex function in its second argument, $\Gamma[h^{(n)},u]$ is a positive semi-definite matrix. Therefore, we obtain
$J_{T}[h^{(n)},u]=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]\geq I>0$
This concludes the proof.
Appendix E Proof of Theorem 4
Let $V[\cdot]$, $W[\cdot]$ be the functionals defined in (23) and (24), respectively. Moreover, let $\hat{W}[\cdot]$ be a functional obtained by the alternating sequence $\hat{h}^{(n)}\to\hat{Q}_{Y|X}^{(n+1)}\to\hat{h}^{(n+1)}$ substituting $\hat{Q}_{Y|X}$ with fixed $\hat{h}$ as follows
$\begin{split}\hat{W}[\hat{h}^{(n)}]=&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}\hat{h}^{(n)}(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}\end{split}$ (29)
where $A^{(n)}=A[\hat{h}^{(n)}]$. Similarly to Theorem 2, we let $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and consider that
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{h}^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}\hat{h}^{(n)}(i)A^{(n)}(x,i,s)\right).\end{split}$
Substituting the definition of $A(x,,s)$ and using (26), we obtain
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\leq\quad V[h^{*}]-\hat{W}[\hat{h}^{(n)}]\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}(p_{X}||h^{*})+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad\qquad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}.\end{split}$ (30)
Note that the right-hand side of (30) can be bounded by
$\begin{split}&\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=D_{f}\left(p_{X}||v^{(n)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{i\in\mathcal{X}}(h^{*}-v^{(n)})\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &\quad+D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||v^{(n)}\right)+\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &=-\left[D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),v^{(n)}}(h^{*})\right]+\left[D_{f}(p_{X}||h^{(n+1)})-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right]\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\end{split}$
where $T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)$ is the first order Taylor expansion of $D_{f}(p_{X}||\cdot)$, centered in $v^{(n)}$ and evaluated in $h^{(n+1)}$, and $(a)$ is verified since $D_{f}(p_{X}||\cdot)$ is a convex function in its second argument.
Since $\hat{W}[\hat{h}^{(n)}]\geq W[\hat{h}^{(n)}]\geq V[h^{*}]$ we can rewrite
$\begin{split}0\leq\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq G[\hat{h}^{(n)}]\end{split}$
where
$\begin{split}G[\hat{h}^{(n)}]=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)+s_{P}\left[D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right].\end{split}$
Summing over $N$ terms we obtain
$\begin{split}\sum_{n=1}^{N}\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)\\ &+s_{P}\sum_{n=1}^{N}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\\ \leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{\hat{Q}_{Y|X}^{(1)}}\right)+\sum_{n=1}^{N}o\left(||h^{(n+1)}-v^{(n)}||\right)\\ \leq&\quad\tilde{L}(h^{*},h^{(0)})\end{split}$
where $\tilde{L}(h^{*},h^{(0)})$ is finite if the limit $\lim_{n\to\infty}||h^{(n+1)}-v^{(n)}||=0$ converges at least linearly. Thus, we can rewrite
$0\leq\lim_{N\to\infty}\sum_{n=1}^{N}\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq\tilde{L}(h^{*},h^{(0)})$
proving the convergence as in Theorem 2. This completes the proof.
Appendix F Proof of Theorem 5
We start by first introducing the following auxiliary lemma.
Lemma 9.
Let $s_{D}\geq 0$, $s_{P}\in[0,s_{P,\max})$ be given with $s=(s_{D},s_{P})$ and let $Q_{Y|X}$ be a transition matrix included in the set $\mathcal{L}_{(D,P)}$ defined as follows:
$\mathcal{L}_{(D,P)}=\{Q_{Y|X}:\mathbb{E}_{Q_{Y|X}}\left[d(x,y)\right]\leq D\land D_{f}(p_{X}||q_{Y})\leq P\}.$
Furthermore, let $\Lambda_{s,v[q_{Y}]}$ be the set defined as
$\Lambda_{s,v[q_{Y}]}\triangleq\big{\{}\lambda\in\mathbb{R}^{|\mathcal{X}|}:\forall x\in\mathcal{X},\lambda(x)\geq 0\land\forall y\in\mathcal{X},~{}\sum_{x\in\mathcal{X}}p_{X}(x)\lambda(x)A(x,,s)\leq 1\big{\}}$
Then, $\forall\mathbf{\lambda}\in\Lambda_{s,v[q_{Y}]}$, we obtain
$\begin{split}R(D,P)\geq&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)-s_{D}D-s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y).\end{split}$
Proof:
Let $\mathbf{\lambda}\in\Lambda_{s,v[q_{Y}]}$ and $Q_{Y|X}\in\mathcal{L}_{(D,P)}$, then:
$\begin{split}&I\left(p_{X},Q_{Y|X}\right)+\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)+s_{D}D+s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y)\\ &\geq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\log\left(\frac{Q_{Y|X}}{q_{Y}\lambda(x)A(x,,s)}\right)\\ &\geq 1-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\left(\frac{q_{Y}\lambda{x}A(x,,s)}{Q_{Y|X}}\right)\\ &=1-\sum_{y\in\mathcal{X}}q_{Y}\sum_{x\in\mathcal{X}}p_{X}\lambda(x)A(x,,s)\\ &\geq 1-\sum_{y\in\mathcal{X}}q_{Y}=0.\end{split}$
The equality is ensured by the possibility of choosing $\lambda(x)\in\Lambda_{s,v[q_{Y}]}$ as
$\begin{split}\lambda(x)=\frac{1}{\sum_{y\in\mathcal{X}}q_{Y}A(x,,s)}\\ \end{split}$
that once substituted describes the optimization problem found in Corollary 1. This completes the proof of the lemma. ∎
Now we use Lemma 9 to prove Theorem 5. In particular, (18) can be derived from the following inequality:
$\begin{split}R(D,P)&\leq\quad I\left(p_{X},\hat{Q}_{Y|X}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)\hat{Q}_{Y|X}(y|x)\log\left(\frac{\hat{Q}_{Y|X}(y,x)}{\sum_{x\in\mathcal{X}}p_{X}(x)\hat{Q}_{Y|X}(y|x)}\right)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+s_{P}\sum_{y\in\mathcal{X}}q_{Y}(y)\frac{p_{X}(y)}{v[q_{Y}](y)}\partial f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)\\ &\quad-s_{P}\Bigg{\{}\sum_{y\in\mathcal{X}}q_{Y}(y)f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)-P\Bigg{\}}-s_{P}\sum_{j\in\mathcal{X}}\hat{h}(j)A(x,j,s)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+\hat{W}[\hat{h}].\end{split}$
(17) is derived as an application of Lemma 9 by choosing $\lambda(x)$ as
$\lambda(x)=\left(c_{\max}\sum_{y\in\mathcal{X}}\hat{h}(y)A(x,,s)\right)$
which respects the assumption of the theorem. This completes the proof.
Appendix G Proof of Theorem 6
The functional form of $J[h](i,j)$ in the case of Theorem 2 is obtained from
$\begin{split}\frac{\partial S[h](i)}{\partial h(j)}&=h(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{\partial}{\partial h(j)}\left(\frac{A[S[h]](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s)}\right)+c[h,S[h]](i)\delta_{i,j}\end{split}$
where
$\begin{split}\frac{\partial(\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s))}{\partial h(j)}\ =\sum_{k\in\mathcal{X}}h(k)\frac{\partial A[S[h]](x,k,s)}{\partial h(j)}+A[S[h]](x,j,s)\end{split}$
$\begin{split}\frac{\partial A(x,i,s)}{\partial h(j)}=-s_{P}A[S[h]](x,i,s)\frac{p_{X}(i)^{2}}{(S[h(i)])^{3}}\partial f\left(\frac{p_{X}(i)}{S[h(i)]}\right)\frac{\partial S[h](i)}{\partial h(j)}\end{split}.$
By noticing that
$\frac{p_{X}(i)^{2}}{(S[h](i))^{3}}\partial f\left(\frac{p_{X}(k)}{S[h](i)}\right)=\frac{\partial^{2}}{\partial q(i)^{2}}D_{f}(p_{X}||q)\Big{|}_{S[h]}$
and defining the matrices $M$ and $\Gamma$ as in (LABEL:matrix:_Newton_M) and (14), we can rewrite the entries of $J[h^{*}]$ as
$\begin{split}J[h^{*}](i,j)&=c[h^{*},h^{*}](i)\left(\delta_{i,j}-\Gamma[h^{*},h^{*}](i)J[h^{*}](i,j)\right)\\ &\quad+\sum_{k\in\mathcal{X}}\Gamma[h^{*},h^{*}](k)M[h^{*},h^{*}](i,k)J[h^{*}](k,j)-M[h^{*},h^{*}](i,j)\\ \end{split}$ (31)
where $\delta_{i,j}$ is the Kronecker delta. As a result, we can express (31) in matrix form as follows
$\begin{split}J[h^{*}]&=C[h^{*},h^{*}](I-\Gamma[h^{*},h^{*}]J[h^{*}])-M[h^{*},h^{*}]+M[h^{*},h^{*}]\Gamma[h^{*},h^{*}]J[h^{*}]\\ &=(C[h^{*},h^{*}]-M[h^{*},h^{*}])(I-\Gamma J[h^{*}])\end{split}$
where $C[\cdot,\cdot]$ is defined in (15). Finally, we obtain (20) noticing that $C[h^{*},h^{*}]=I$ due to the optimality conditions found in Lemma 2, thus concluding the proof.
Appendix H Proof of Lemma 5
Let matrices $\Phi$ and $Q$ be defined as
$\Phi$ $\triangleq\left[\sqrt{p_{X}(x)}\frac{A[h^{*}](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[h^{*}](x,k,s)}\right]_{(i,x)\in\mathcal{X}^{2}}$
$Q$ $\triangleq\operatorname{diag}\Big{[}h^{*}(i)\Big{]}_{i\in\mathcal{X}}.$
Then, the following identity can be verified
$Q^{\frac{1}{2}}M^{*}Q^{-\frac{1}{2}}=Q^{\frac{1}{2}}\Phi\Phi^{T}Q^{\frac{1}{2}}=(Q^{\frac{1}{2}}\Phi)(Q^{\frac{1}{2}}\Phi)^{T}$
where $Q^{\frac{1}{2}}M^{*}Q^{-\frac{1}{2}}$ is necessarily symmetric and at least semi-positive definite. To guarantee positive definiteness of $Q^{\frac{1}{2}}M^{*}Q^{-\frac{1}{2}}$, and thus the fact that the eigenvalues of $M$ are strictly positive, we need to impose conditions on the full rank of $\Phi$. To address them, we can factorize $\Phi$ into the product $\Phi=UDV$, where
$D$ $=\left[e^{-s_{D}d(i,j)}\right]_{(i,j)\in\mathcal{X}\times\mathcal{X}}$
$U$ $=\operatorname{diag}\left[e^{-s_{P}g(p_{X},h^{*},i)}\right]_{i\in\mathcal{X}}$
$V$ $=\operatorname{diag}\left[\frac{\sqrt{p_{X}(x)}}{\sum_{k\in\mathcal{X}}h(k)A[h^{*}](x,k,s)}\right]_{x\in\mathcal{X}}.$
Since both $U$ and $V$ are positive definite matrices, it is easy to verify that $\Phi$ is a full-rank matrix if and only if $D$ is full rank too. This completes the proof.
Appendix I Proof of Lemma 6
We can verify that the lemma is the result of the Gershgorin Circle Theorem [45] applied to the columns of $M^{*}$. Noticing that all entries $M^{*}(i,j)$ are strictly positive, the disk radius $R(j)$ for column $j$ is:
$R(j)+M^{*}(i,i)=\sum_{i\in\mathcal{X}}M^{*}(i,j)$
where
$\begin{split}\sum_{i\in\mathcal{X}}M^{*}(i,j)&=\sum_{i\in\mathcal{X}}h^{*}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,i,s)A[h^{*}](x,j,s){}}{(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s))^{2}}\\ &=\sum_{x\in\mathcal{X}}\frac{A[h^{*}](x,j,s)}{\left(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s)\right)^{2}}\left(\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)\right)\\ &=c[h^{*},h^{*}](j)=1.\end{split}$
Thus the eigenvalues of $M^{*}$ are each in at least one of the disks $I_{i}=\{z\in\mathbb{R}:|z-M^{*}(i,i)|\leq 1-M^{*}(i,i)\}$, which are all contained in the disk $I=\{z\in\mathbb{R}:|z|\leq 1\}$. This completes the proof.
Appendix J Proof of Theorem 7
Using Lemmas 5 and 6, the following inequalities hold
$1\displaystyle 0<\operatorname{eig}(M^{*})\leq 1\implies 0\leq\operatorname{eig}(I-M^{*})<$ 1
$0\leq\operatorname{eig}(I-M^{*})<1\xRightarrow{(a)}1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)$
$1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)\implies 0<\operatorname{eig}((I+(I-M^{*})\Gamma^{*})^{-1})\leq 1$
where $(a)$ is due to $\Gamma$ being a positive definite matrix. Using the previous inequalities, we can rewrite (20) as follows:
$J(h^{*})=(I+(I-M^{*})\Gamma^{*})^{-1}(I-M^{*}).$
Define $\theta_{\sup}\triangleq\operatorname{eig}_{\max}(I-M^{*})\cdot\operatorname{eig}_{\max}((I+(I-M^{*})\Gamma^{*})^{-1})$. Then, we can show that $10\leq\operatorname{eig}(J[h^{*}])\leq\theta_{\sup}<$ 1 is always verified. The second part of the theorem follows directly from [42, Theorem 5] hence we omit it. This completes the proof.
Appendix K Proof of Theorem 8
Since $R(D,P)$ is a non-increasing convex function in $D_{s}$ and $P_{s}$, we can derive:
$\frac{\partial^{2}R(D,P)}{\partial P_{s}^{2}}=-\frac{\partial s_{P}}{\partial P_{s}}\geq 0$
meaning that more constrained values of $P_{s}$ are associated with larger $s_{P}$. Thus, for a given $s_{D}$, let $s_{P,\max}$ be the value of the Lagrangian $s_{P}$ associated with the constraint $P=0$. Then the solution $h^{*}$ is necessarily unique and must be $h^{*}=p_{X}$. Then, due to the properties of the Jacobian $J(h^{*})$,
$\begin{split}(I-\Gamma J(h^{*}))\geq 0\implies s_{P,\max}\leq\frac{1}{\theta_{\max}f^{\prime\prime}(0)}.\end{split}$
In order to guarantee $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1, it is sufficient to have
$\begin{split}(I-\Gamma)\geq 0\implies s_{P,\max}\leq\frac{1}{f^{\prime\prime}(0)}.\end{split}$
Since in the non-degenerate case we have $10<\theta_{\max}<$ 1, we can construct $\epsilon^{\prime}=\min\left\{s_{P,\max},\frac{1}{f^{\prime\prime}(0)}\left(\frac{1}{\theta_{\max}}-1\right)\right\}$ > 0 and define the set $I_{s_{P}}^{\epsilon^{\prime}}$ where the exponential convergence of the approximate algorithm is guaranteed. This concludes the proof.
Appendix L Proof of Lemma 8
We remind that $f=|x-1|$ is the function associated with the TV distance $TV(\cdot||\cdot)$. The first statement can be proved by first establishing the uniform convergence of $f_{n}\to f$ as $n\to\infty$
$\sup_{x\in\mathbb{R}}\left|f_{n}(x)-f(x)\right|$ $=\sup_{x\in\mathbb{R}}\left|\frac{2}{\pi}(x-1)\arctan(n(x-1))-|x-1|\right|$
$=\sup_{x\in\mathbb{R}^{+}_{0}}\left|\frac{2}{\pi}x\arctan(nx)-x\right|$
$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\left(\frac{\pi}{2}-\arctan(nx)\right)$
$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\arctan\left(\frac{1}{nx}\right)$
$\leq\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{n\pi}=\frac{2}{n\pi}$
meaning that $\lim_{n\to\infty}\sup_{x\in\mathbb{R}}|f_{n}(x)-f(x)|=0$. A direct consequence of the above is that $D_{f_{n}}\to TV$ uniformly in the limit of $n\to\infty$ since, for any $p,q\in\mathcal{P}(\mathcal{X})$,
$\lim_{n\to\infty}|D_{f_{n}}(p||q)-TV(p||q)|$ $=\lim_{n\to\infty}\left|\sum_{x\in\mathcal{X}}q(x)\left(f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right)\right|$
$\leq\sum_{x\in\mathcal{X}}q(x)\lim_{n\to\infty}\left|f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right|=0.$
Instead, the inequality $f_{n}(x)\leq f(x),\forall x\in\mathbb{R}$, implies that for all $n\in\mathbb{N}$ and $\forall p,q\in\mathcal{P}(\mathcal{X})$, the inequality $D_{f_{n}}(p||q)\leq TV(p||q)$ holds. This concludes the proof.

### v2

@@
 Abstract
-We study the computation of the rate-distortion-perception function (RDPF) for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences.
-In this setting, the RDPF forms a convex programming problem for which we characterize the optimal parametric solutions.
+We study the computation of the rate-distortion-perception function (RDPF) for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint belonging to the family of $f$-divergences.
+In this setting, the RDPF forms a convex programming problem for which we characterize optimal parametric solutions.
 We employ the developed solutions in an alternating minimization scheme, namely Optimal Alternating Minimization (OAM), for which we provide convergence guarantees.
-Nevertheless, the OAM scheme does not lead to a direct implementation of a generalized Blahut-Arimoto (BA) type of algorithm due to the presence of implicit equations in the structure of the iteration.
-To overcome this difficulty, we propose two alternative minimization approaches whose applicability depends on the smoothness of the used perception metric: a Newton-based Alternating Minimization (NAM) scheme, relying on Newton’s root-finding method for the approximation of the optimal iteration solution, and a Relaxed Alternating Minimization (RAM) scheme, based on a relaxation of the OAM iterates.
-Both schemes are shown, via the derivation of necessary and sufficient conditions, to guarantee convergence to a globally optimal solution.
-We also provide sufficient conditions on the distortion and the perception constraints which guarantee that the proposed algorithms converge exponentially fast in the number of iteration steps.
-We corroborate our theoretical results with numerical simulations and draw connections with existing results.
+Nevertheless, the OAM scheme does not lead to a direct implementation of a generalized Blahut-Arimoto (BA) type of algorithm due to implicit equations in the iteration’s structure.
+To overcome this difficulty, we propose two alternative minimization approaches whose applicability depends on the smoothness of the used perception metric: a Newton-based Alternating Minimization (NAM) scheme, relying on Newton’s root-finding method for the approximation of the optimal solution of the iteration, and a Relaxed Alternating Minimization (RAM) scheme, based on relaxing the OAM iterates.
+We show, by deriving necessary and sufficient conditions, that both schemes guarantee convergence to a globally optimal solution.
+We also provide sufficient conditions on the distortion and perception constraints, which guarantee that the proposed algorithms converge exponentially fast in the number of iteration steps.
+We corroborate our theoretical results with numerical simulations and establish connections with existing results.
+Index Terms:
+Lossy Source Coding, Rate-Distortion-Perception Trade-off, Blahut-Arimoto Algorithms, Approximation Methods, $f$-divergence
 I Introduction
 The rate-distortion-perception (RDP) trade-off studies the relevant problem of lossy compression under perceptual constraints on the reconstructed samples, generalizing the classical rate-distortion (RD) analysis [2].
-Concurrently proposed by Blau and Michaeli in [3] and Matsumoto in [4, 5], its introduction is motivated by the increasing necessity of a theoretical framework able to incorporate observations raised by a wide body of research spanning from machine learning to multimedia applications, see e.g., [6, 7, 8, 9], remarking the presence of an inherent tension between perceptual quality and fidelity of the compressed samples.
-In this context, perceptual quality refers to the property of a sample to appear pleasing from the human standpoint.
-Empirical evaluations of perceptual quality using human scoring [10, 11, 12] show that conventional fidelity measures fail to capture human preference and perception, especially in extreme compression scenarios.
-Hence, classical rate-distortion theory, focusing solely on the fidelity of the reconstructed samples, may not provide an adequate theoretical framework when considering “human-oriented” data sources, such as images, audio, and video.
+Concurrently proposed by Blau and Michaeli in [3] and Matsumoto in [4, 5], its introduction is motivated by the increasing necessity of a theoretical framework able to incorporate observations raised by a wide body of research spanning from machine learning to multimedia applications; see e.g., [6, 7, 8, 9], highlighting the presence of an inherent tension between perceptual quality and fidelity of the compressed samples.
+In this context, perceptual quality refers to the property of a sample to appear pleasing from a human standpoint.
+Empirical evaluations of perceptual quality using human scoring [10, 11, 12] show that conventional fidelity measures fail to capture human preferences and perceptions, especially in extreme compression scenarios.
+Hence, classical rate-distortion theory, focusing solely on the fidelity of the reconstructed samples, may not provide an adequate theoretical framework when applied to “human-oriented” data sources, such as images, audio, and video.
 The mathematical representation of the RDP trade-off is embodied by the rate-distortion-perception function (RDPF), which complements the classical single-letter rate-distortion function (RDF) with a divergence constraint between the source and reconstruction distributions.
-The additional constraint acts as a proxy of human perception, measuring the deviation from the real source distribution, also referred to as “natural scene statistic”, following similar principles behind a class of no-reference image quality metrics [13, 14].
-It is worth noticing, however, that the selection of specific divergence metrics may be application-dependent and is still an active area of research.
-An alternative interpretation of the divergence constraints can be found in its potential as a semantic quality metric, i.e.
-a quantification of the importance of the reconstructed source from the observer’s perspective [15].
-For example, in [16] a comparison between the segmentation capabilities of models trained on traditionally compressed samples against compressed samples with enhanced perceptual quality through Generative Adversarial Network (GAN)-based restoration shows a remarkable improvement in segmentation performances, especially for smaller scene objects.
+The additional constraint acts as a proxy for human perception, measuring the deviation from the real source distribution, also referred to as “natural scene statistic”, following similar principles to those behind a class of no-reference image quality metrics [13, 14].
+However, it is worth noting that the selection of specific divergence metrics may be application-dependent and is still an active area of research.
+An alternative interpretation of the divergence constraints can be found in its potential as a semantic quality metric, i.e., a quantification of the importance of the reconstructed source to the observer [15].
+For example, in [16], a comparison between the segmentation capabilities of models trained on traditionally compressed samples against compressed samples with enhanced perceptual quality through Generative Adversarial Network (GAN)-based restoration shows a remarkable improvement in segmentation performance, especially for smaller scene objects.
 I-A Related Work
-Since its introduction, the RDP trade-off received substantial interest from the information theory community, deriving operational characterizations in a variety of operational scenarios.
-Focusing on the case where infinite common randomness is available both at the encoder and the decoder, Theis and Wagner [17] provide variable-length codes for both the one-shot and asymptotic regime exploiting the properties of the strong functional representation lemma [18].
+Since its introduction, the RDP trade-off has received substantial interest from the information theory community, deriving operational characterizations in a variety of operational scenarios.
+Focusing on the case where infinite common randomness is available both at the encoder and the decoder, Theis and Wagner [17] provide variable-length codes for both the one-shot and asymptotic regime, exploiting the properties of the strong functional representation lemma [18].
 In the context of the output-constrained RDF, but also valid for the "perfect realism" RDPF case, Saldi et al.
@@
 In [20], Chen et al.
-derives coding theorems for the asymptotic case, focusing on the differences between three operational cases defined by the availability of randomness between encoder and decoder, i.e.
-infinite common randomness, only private randomness, and the deterministic case.
+derive coding theorems for the asymptotic case, focusing on the differences between three operational cases defined by the availability of randomness between encoder and decoder, i.e., infinite common randomness, only private randomness, and the deterministic case.
 Similar to its classical counterpart, the RDPF does not enjoy any general analytical solution.
-However, for specific source distributions and particular choices of the distortion and perception measures, there exist closed-form expressions, such as for binary sources subject to Hamming distortion and total variation (TV) distance [3] and Gaussian sources under mean squared-error distortion and various perception measures [21].
-The available closed forms, while providing theoretical insights on the RDP trade-off, have limited applicability when considering arbitrary distortion and perception metrics.
+However, for specific source distributions and particular choices of distortion and perception measures, there exist closed-form expressions, such as for binary sources subject to Hamming distortion and total variation (TV) distance [3] and Gaussian sources under mean squared-error distortion and various perception measures [21, 22, 23].
+The available closed forms, while providing theoretical insights into the RDP trade-off, have limited applicability when considering arbitrary distortion and perception metrics.
 In the case of the classical RDF, the introduction of the celebrated Blahut-Arimoto (BA) algorithm [24] mitigates the absence of a general closed-form solution, allowing for the exact computation of the RDF for the discrete case.
-Furthermore, its versatility inspired various approximation methods for the RD in the continuous case [23, 24] and allowed it to be adapted to a variety of source coding problems, with applications from quantum information theory [25] to goal-oriented communication [26].
-Also in the RDPF case, numerical solutions have been explored to some extent.
-Data-driven solutions have been proposed, usually employing a generative adversarial scheme minimizing a linear combination of distortion and perception metrics, see e.g., [3, 27, 28, 29].
-Despite the advantage of directly optimizing the image/video codec using only source samples, these methodologies still require particular effort as they are generally highly computational and data-intensive and may suffer from a lack of generalization capabilities.
-Algorithmic solutions for the RDPF estimation are also present.
+Furthermore, its versatility inspired various approximation methods for the RDF in the continuous case [25, 26] and enabled its adaptation to a variety of source coding problems, with applications from quantum information theory [27] to goal-oriented communication [28].
+Moreover, in the RDPF case, numerical solutions have been explored to some extent.
+Data-driven solutions have been proposed, usually employing a generative adversarial scheme minimizing a linear combination of distortion and perception metrics, see, e.g., [3, 22, 29, 30].
+Despite the advantage of directly optimizing the image/video codec using only source samples, these methodologies still require considerable effort, as they are generally highly computationally demanding and data-intensive and may suffer from a lack of generalization capabilities.
+Algorithmic solutions for estimating the RDPF also exist.
 In the case of discrete alphabets, Chen et al.
-in [30] cast the RDPF problem as an entropic-regularized Wasserstein barycenter problem and propose a solving method based on the Sinkhorn algorithm applicable to arbitrary distortion measures and perception measure being either a Wasserstein-type distance, the Kullback–Leibler divergence, or the TV distance.
+in [31] cast the RDPF problem as an entropic-regularized Wasserstein barycenter problem and propose a solution method based on the Sinkhorn algorithm applicable to arbitrary distortion measures and with the perception measure being either a Wasserstein-type distance, the Kullback–Leibler divergence, or the TV distance.
 Focusing on Gaussian sources, Serra et al.
-in [21] designs an alternating minimization method for the computation of the Gaussian RDPF for arbitrary fidelity and perception measures, deriving analytical solutions for the so-called “perfect realism” case.
-Always for the “perfect realism” case, Serra et al.
-[31] design a solving algorithm for general multivariate continuous sources and distortion metrics, leveraging the information-geometric aspects of the constrained rate-distortion problem.
-To the best of our knowledge, none of the existing methodologies addresses the generic computation of the RDPF for discrete sources relying on generalizations of the classical BA algorithm.
+in [21] design an alternating minimization method for the computation of the Gaussian RDPF for arbitrary fidelity and perception measures, deriving analytical solutions for the so-called “perfect realism” case.
+Furthermore, for the “perfect realism” case, Serra et al.
+[32] design a solution algorithm for general multivariate continuous sources and distortion metrics, leveraging the information-geometric aspects of the constrained rate-distortion problem.
+To the best of our knowledge, none of the existing methodologies address the generic computation of the RDPF for discrete sources relying on generalizations of the classical BA algorithm.
 I-B Contributions
-The objective of this work is to propose a generic algorithmic approach for the computation of the RDPF, focusing on the case of discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the class of $f$-divergences.
-Our results leverage the fact that the RDPF forms a convex program under mild regularity conditions on the perception constraint (i.e., convexity on the second argument), which are satisfied by the considered class of divergences.
-This enables us to derive a parametric characterization of the optimal solution of the RDPF (Lemma 1), which is subsequently utilized to construct an alternating minimization procedure, thereon referred to as the Optimal Alternating Minimization (OAM) scheme, for which we also establish convergence guarantees (Theorem 2).
-However, the resulting structure of the OAM scheme relies on a set of implicit equations in the variables of interest, thus prohibiting the direct implementation of a generic BA algorithm similar to what is already known for the classical rate-distortion theory for $\operatorname{i.i.d.}$ sources and single-letter distortions [22].
-Motivated by this technical difficulty, we propose two alternative minimization approaches that solve the implementability issue and whose applicability depends on the smoothness of the considered perception function.
-  • In Section III-A, we design a Newton-based Alternating Minimization (NAM) scheme where, noticing that the solution of the OAM iterate is equivalent to a root-finding problem (Lemma 3), we use Newton’s root-finding method [32] for the computation of the optimal iteration step (Theorem 3).
-  • In Section III-B, we introduce a Relaxed Alternating Minimization (RAM) scheme, where we leverage a new relaxed formulation of the structure of the iterations in the OAM scheme and, subsequently, we derive necessary and sufficient conditions to ensure convergence to a globally optimal solution (Theorem 4).
-In Section IV we design the algorithmic embodiment for the NAM and RAM schemes (see Algorithm 1 and Algorithm 2, respectively) and develop suitable stopping criteria for both algorithms (Theorem 5).
-Moreover, we also provide sufficient conditions on the structure of the distortion and the perception constraints based on which our algorithms converge exponentially fast in the number of iterations (Theorems 7 and 8).
-We corroborate our theoretical findings with numerical simulations (Section V), with emphasis on the case of the TV perception metric, for which we develop a smooth approximation (Lemma 8).
+The objective of this work is to propose a generic algorithmic approach for the computation of the RDPF, focusing on the case of discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint belonging to the class of $f$-divergences.
+Our results leverage the fact that the RDPF forms a convex program under mild regularity conditions on the perception constraint (specifically, convexity in the second argument), which are satisfied by the considered class of divergences.
+This enables us to derive a parametric characterization of the optimal solution of the RDPF (Lemma 1), which is subsequently utilized to construct an alternating minimization procedure, hereafter referred to as the Optimal Alternating Minimization (OAM) scheme, for which we also establish convergence guarantees (Theorem 2).
+However, the resulting structure of the OAM scheme relies on a set of implicit equations in the variables of interest, thus preventing the direct implementation of a generic BA algorithm, as is already known for the classical rate-distortion theory for $\operatorname{i.i.d.}$ sources and single-letter distortions [24].
+Motivated by this technical difficulty, we propose two alternative minimization approaches that address the implementability issue and whose applicability depends on the smoothness of the considered perception function.
+  • In Section III-A, we design a Newton-based Alternating Minimization (NAM) scheme observing that the solution of the OAM iterate is equivalent to a root-finding problem (Lemma 3), which allows us to apply Newton’s root-finding method [33] to compute the optimal iteration step (Theorem 3).
+  • In Section III-B, we introduce a Relaxed Alternating Minimization (RAM) scheme, where we leverage a new relaxed formulation of the structure of the OAM iterations and subsequently, we derive necessary and sufficient conditions to ensure convergence to a globally optimal solution (Theorem 4).
+In Section IV, we design the algorithmic implementations of the NAM and RAM schemes (see Algorithm 1 and Algorithm 2, respectively) and develop suitable stopping criteria for both algorithms (Theorem 5).
+Moreover, we also provide sufficient conditions on the structure of the distortion and perception constraints under which our algorithms converge exponentially fast in the number of iterations (Theorems 7 and 8).
+We corroborate our theoretical findings with numerical simulations (Section V), with emphasis on the TV perception metric, for which we develop a smooth approximation (Lemma 8).
 I-C Notation
-Let $\mathbb{N}$ denote the set of natural numbers, $[a:b]\subset\mathbb{N}$ the integer interval including its extremes, $\mathbb{R}$ the set of real numbers, and $\mathbb{R}_{0}^{+}$ the set of non-negative real numbers.
+Let $\mathbb{N}$ denote the set of natural numbers, $a:b]\subset\mathbb{N}$ the integer interval including its extremes, $\mathbb{R}$ the set of real numbers, and $\mathbb{R}_{0}^{+}$ the set of non-negative real numbers.
 Given a discrete alphabet $\mathcal{X}$, we denote by $(\mathcal{X},\mathbb{B}(\mathcal{X}))$ the Borel measurable space induced by the metric, with $\mathcal{P}(\mathcal{X})$ denoting the set of probability measures defined thereon.
@@
 We denote with $C^{n}$ the set of $n^{th}$-times differentiable functions.
-Given a function $f\in C^{0}$, we denote with $\partial f$ its sub-gradient [33, Definition 8.3], while, if $f\in C^{2}$, we denote by $f^{\prime\prime}(\cdot)$ the second derivative with respect to its argument.
-We denote by $D(\cdot||\cdot)$ a generic divergence measure whereas $D_{f}(\cdot||\cdot)$ denotes a divergence belonging to the class of $f$-divergences.
+Given a function $f\in C^{0}$, we denote with $\partial f$ its sub-gradient [34, Definition 8.3], while, if $f\in C^{2}$, we denote by $f^{\prime\prime}(\cdot)$ the second derivative with respect to its argument.
+We denote by $D(\cdot||\cdot)$ a generic divergence measure, whereas $D_{f}(\cdot||\cdot)$ denotes a divergence belonging to the class of $f$-divergences.
 Given a vector $v\in\mathbb{R}^{d}$, we indicate with $\operatorname{diag}{v}\in\mathbb{R}^{d\times d}$ the matrix with as diagonal the elements of $v$ and zeros otherwise.
@@
 II Preliminaries
-We start this section by providing the formal definition of the RDPF and an overview of its operational meaning, following the works of [3, 17].
-Subsequently, given their relevance in this work, we introduce the category of statistical divergences known as $f$-divergences and conclude by providing an overview of the alternating minimization methodology.
+We start this section by providing the formal definition of the RDPF and an overview of its operational meaning, following [3, 17].
+Subsequently, given their relevance to this work, we introduce the category of statistical divergences known as $f$-divergences and conclude with an overview of the alternating minimization methodology.
 II-A Rate-Distortion-Perception Functions
-We consider finite alphabet sources and stochastic encoders/decoders pairs having access to a common source of randomness and define the minimum achievable rates when we consider per-letter expected distortion and per-letter perception constraints.
+We consider finite alphabet sources and stochastic encoder/decoder pairs having access to a common source of randomness and define the minimum achievable rates under per-letter expected distortion and per-letter perception constraints.
 We assume that we are given an $\operatorname{i.i.d.}$ sequence of $n$-length random variables ${X}^{n}\in\mathcal{X}^{n}$ that induce the probability distribution $p_{X}\in\mathcal{P}(\mathcal{X})$.
-Formally, a stochastic encoder $f^{n}_{E}$ is any function belonging to the set $\mathcal{F}^{n}_{E}=\{f:\mathcal{X}^{n}\times\mathbb{R}\to\mathbb{N}\}$, whereas a stochastic decoder $g^{n}_{D}$ is any function in the set $\mathcal{G}^{n}_{D}=\{g:\mathbb{N}\times\mathbb{R}\to\mathcal{X}^{n}\}$.
+Formally, a stochastic encoder $f^{n}_{E}$ is any function in the set $\mathcal{F}^{n}_{E}=\{f:\mathcal{X}^{n}\times\mathbb{R}\to\mathbb{N}\}$, whereas a stochastic decoder $g^{n}_{D}$ is any function in the set $\mathcal{G}^{n}_{D}=\{g:\mathbb{N}\times\mathbb{R}\to\mathcal{X}^{n}\}$.
 A stochastic code is an element of $\mathcal{F}^{n}_{E}\times\mathcal{G}^{n}_{D}$.
-Without loss of generality, the randomness at the encoder and decoder is represented as a single real number, i.e., an infinite number of bits, and is assumed shared by the pair, i.e., common randomness.
-We let $d:\mathcal{X}^{2}\to\mathbb{R}_{0}^{+}$ denote a single-letter distortion function and $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ denote a divergence measure.
+Without loss of generality, the randomness at the encoder and decoder is modeled as a single real number (i.e., representing an infinite number of bits), and is assumed shared by the pair, i.e., common randomness.
+We let $d:\mathcal{X}^{2}\to\mathbb{R}_{0}^{+}$ denote a single-letter distortion function and $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ denote a divergence function.
 Moreover, we define the sets of fidelity criteria $\{\Delta_{i}\}_{i\in[1:n]}$ and $\{\Phi_{i}\}_{i\in[1:n]}$ as follows
 $\Delta_{i}\triangleq\mathbb{E}_{p_{X_{i},Y_{i}}}\left[d(X_{i},Y_{i})\right],\qquad\Phi_{i}\triangleq D(p_{X_{i}}||q_{Y_{i}})$
-where $\Delta_{i}$ is the expected $i^{th}$ symbol distortion and $\Phi_{i}$ is the $i^{th}$ symbol divergence with respect to the reconstructed symbol $Y_{i}$.
+where $\Delta_{i}$ is the expected distortion of the $i^{th}$ symbol and $\Phi_{i}$ is the $i^{th}$ symbol divergence with respect to the reconstructed symbol $Y_{i}$.
 We are now ready to introduce the definition of achievability and that of the infimum of all achievable rates.
@@
 Then, we define
-[…]
-Next, we give the definition of the information-theoretic characterization of the RDPF [3].
+$R_{cr}(D,P)\triangleq\inf\{R:R\text{ is $(D,P)$-achievable}\}.$
+Next, we state the definition of the information-theoretic characterization of the RDPF [3].
 Definition 2.
@@
 $\quad D(p_{X}||q_{Y})\leq P$ (3)
-where $D\in[D_{\min},D_{\max}]\subseteq[0,\infty)$, $P\in[P_{\min},P_{\max}]\subseteq[0,\infty)$,
-$q_{Y}=\sum_{x\in\mathcal{X}}Q_{Y|X}p_{X}\qquad\text{and}\qquad I(X,Y)=D_{KL}(p_{X}Q_{Y|X}||p_{X}q_{Y})\triangleq I(p_{X},Q_{Y|X})$
+where $D\in[D_{\min},D_{\max}]\subseteq[0,\infty)$, $P\in[P_{\min},P_{\max}]\subseteq[0,\infty)$, $q_{Y}=\sum_{x\in\mathcal{X}}Q_{Y|X}p_{X}$, and
+$I(X,Y)=D_{KL}(p_{X}Q_{Y|X}||p_{X}q_{Y})\triangleq I(p_{X},Q_{Y|X})$
 where $I(p_{X},Q_{Y|X})$ highlights the dependency on $\{p_{X},Q_{Y|X}\}$.
-We stress the following technical remarks on Definition 2.
+In what follows, we highlight certain functional properties of Definition 2.
 Remark 1.
-(On Definition 2) Following [3], it can be shown that (1) has some useful properties, under mild regularity conditions.
-In particular, [3, Theorem 1] showed that $R(D,P)$ is (i) monotonically non-increasing function in both $D\in[D_{\min},D_{\max}]\subset[0,\infty)$ and $P\in[P_{\min},P_{\max}]\subset[0,\infty)$; (ii) convex if the divergence $D(\cdot||\cdot)$ is convex in its second argument.
-In the sequel, we assume that in (1), the perception constraint is an $f$-divergence, i.e., $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$, which is known to be convex in both arguments [34, Lemma 4.1].
-Hence in light of the discussion in Remark 1, $R(D,P)$ forms a convex programming problem.
+(On Definition 2 - Functional properties in $(D,P)$) Following [3], it can be shown that (1) has some useful properties, under mild regularity conditions.
+In particular, [3, Theorem 1] showed that, for $D\in[D_{\min},D_{\max}]\subset[0,\infty)$ and $P\in[P_{\min},P_{\max}]\subset[0,\infty)$, $R(D,P)$ is (i) monotonically non-increasing in both $D$ and $P$; (ii) convex in both $D$ and $P$ if the divergence $D(\cdot||\cdot)$ is convex in its second argument.
+Remark 2.
+(On Definition 2 - Functional properties in $Q_{Y|X}$) The program defined by (1)-(3) is convex in the transition matrix $Q_{Y|X}$ for a given $p_{X}$ if the divergence $D(\cdot||\cdot)$ is convex in its second argument, since (1) and (2) are respectively convex and affine functions in $Q_{Y|X}$ [35].
+Furthermore, the identity kernel $Q_{Y|X}=Id$ always satisfies the constraints given by (2)-(3).
+In the sequel, we assume that in (1), the perception constraint is an $f$-divergence, i.e., $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$, which is known to be convex in both arguments [36, Lemma 4.1].
 We conclude this section by providing a theorem that connects $R_{cr}(D,P)$ with $R(D,P)$ for general alphabets.
@@
 under the assumption that
-[…]
-Many commonly used divergences functions belong to the class of $f$-divergences. For example,
+$\text{(i) }f(0)$ $=\lim_{x\to 0^{+}}f(0),\qquad\text{(ii) }0f\left(\frac{0}{0}\right)=0,$
+$\text{(iii) }\forall a\geq 0,~0f\left(\frac{a}{0}\right)=af^{\prime}(\infty).$
+Many commonly used divergence functions belong to the class of $f$-divergences. For example,
   • KL divergence $D_{KL}(\cdot||\cdot)$, obtained by considering $f(x)=x\log(x)$,
@@
   • $\alpha$-divergence $D_{\alpha}(\cdot||\cdot)$, where $f_{\alpha}$ is parameterized by $\alpha\in\mathbb{R}$,
-    $D_{\alpha}(p||q)=\sum_{x\in\mathcal{X}}q(x)f_{\alpha}\left(\frac{p(x)}{q(x)}\right)\qquad$ $f_{\alpha}(x)=\ \begin{cases}\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(\alpha-1)}\qquad\text{if }\alpha\neq 0,\alpha\neq 1\\ x\log(x)-x+1\qquad\text{if }\alpha=1\\ -\ln(x)+x-1\qquad\text{if }\alpha=0\\ \end{cases}.$
-We now provide some general properties of this family of divergences. For any $f$-divergence $D_{f}(\cdot||\cdot)$, the following properties hold:
+    $D_{\alpha}(p||q)$ $=\sum_{x\in\mathcal{X}}q(x)f_{\alpha}\left(\frac{p(x)}{q(x)}\right)\qquad$
+    $f_{\alpha}(x)$ $=\ \begin{cases}\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(\alpha-1)}\qquad\text{if }\alpha\neq 0,\alpha\neq 1\\ x\log(x)-x+1\qquad\text{if }\alpha=1\\ -\ln(x)+x-1\qquad\text{if }\alpha=0\\ \end{cases}.$
+We now state some general properties of this family of divergences. For any $f$-divergence $D_{f}(\cdot||\cdot)$, the following properties hold:
   • (Linearity) $D_{f_{1}+f_{2}}(\cdot||\cdot)=D_{f_{1}}(\cdot||\cdot)+D_{f_{2}}(\cdot||\cdot)$
   • (Joint Convexity) for any $t\in[0,1]$ and $p_{1},p_{2},q_{1},q_{2}\in\mathcal{P}$,
-    $D_{f}(tp_{1}+(1-t)p_{2}||tq_{1}+(1-t)q_{2})\leq tD_{f}(p_{1}||q_{1})+(1-t)D_{f}(p_{2}||q_{2}).$
-  • (Invariance) Let $\hat{f}(x)=f(x)+c(x-1)$ for $c\in\mathbb{R}$, then $D_{f}(\cdot||\cdot)=D_{\hat{f}}(\cdot||\cdot)$.
-The characterization of the family of $f$-divergences here provided summarizes the properties useful for the scope of this work. For a more in-depth mathematical analysis, we refer the reader to [36].
+    $D_{f}(tp_{1}+(1-t)p_{2}||tq_{1}+(1-t)q_{2})$
+    $\qquad\qquad\leq tD_{f}(p_{1}||q_{1})+(1-t)D_{f}(p_{2}||q_{2}).$
+[…]
+The characterization of the family of $f$-divergences provided here summarizes the properties useful for the scope of this work. For a more in-depth mathematical analysis, we refer the reader to [38].
 II-C Alternating Minimization and BA-type algorithms
@@
 $\min_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)$
-where $\mathcal{X}$ and $\mathcal{Y}$ are two arbitrary non-empty sets and the function $f(x,y)$ satisfies $-\infty<f(x,y)\leq+\infty$ for each $x\in\mathcal{X}$ and $y\in\mathcal{Y}$.
+where $\mathcal{X}$ and $\mathcal{Y}$ are two arbitrary non-empty sets and the function $f(x,y)$ satisfies $\infty<f(x,y)\leq+\infty$ for each $x\in\mathcal{X}$ and $y\in\mathcal{Y}$.
 Furthermore, we assume that, for each $x\in\mathcal{X}$, there exists $y\in\mathcal{Y}$ with $f(x,y)<+\infty$, implying that $s\coloneqq\displaystyle\inf_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)<+\infty$.
 Depending on the case, the existence or uniqueness of the minimizer $(x^{*},y^{*})$ such that $f(x^{*},y^{*})=s$ may also be assumed.
-The goal of the alternating minimization method is to create a sequence $\{(x^{(n)},y^{(n)})\}$ such that $\lim_{n\to\infty}f(x^{(n)},y^{(n)})=s$.
-Under specific conditions, such sequence can be defined using the solutions of two sub-problems; for $x_{i}\in\mathcal{X}$, $h(x_{i})=\displaystyle\operatorname*{arg\,min}_{y\in\mathcal{Y}}f(x_{i},y)$ and, for $y_{i}\in\mathcal{Y}$, $g(y_{i})=\displaystyle\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x,y_{i})$.
-For some initial point $y^{(0)}$, we can define the $n$-th element of the sequence as:
-$x^{(n)}=g(y^{(n-1)})\qquad y^{(n)}=g(x^{(n)})\qquad\text{for }n=1,2\ldots.$
+The goal of the alternating minimization method is to construct a sequence $\{(x^{(n)},y^{(n)})\}$ such that $\lim_{n\to\infty}f(x^{(n)},y^{(n)})=s$.
+Under specific conditions, such a sequence can be defined using the solutions of two sub-problems; for $x_{i}\in\mathcal{X}$, $h(x_{i})=\displaystyle\operatorname*{arg\,min}_{y\in\mathcal{Y}}f(x_{i},y)$ and, for $y_{i}\in\mathcal{Y}$, $g(y_{i})=\displaystyle\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x,y_{i})$.
+Starting from an initial point $y^{(0)}$, we can define the $n$-th element of the sequence as:
+$x^{(n)}=g(y^{(n-1)})\qquad y^{(n)}=h(x^{(n)})\qquad\text{for }n=1,2\ldots.$
 Depending on the problem, various sufficient conditions for the existence and optimality of the sequence limit have been studied. For instance,
-  • In [37], Csiszár and Tusnády prove that, if the sequence $\{(x^{(n)},y^{(n)})\}$ guarantees
-    $f(x,y)+f(x,y^{(n-1)})\geq f(x,y^{(n)})+f(x^{(n)},y^{(n-1)})\quad\forall x\in\mathcal{X},\forall y\in\mathcal{Y}$
+  • In [39], Csiszár and Tusnády prove that, if the sequence $\{(x^{(n)},y^{(n)})\}$ guarantees $\forall x\in\mathcal{X},\forall y\in\mathcal{Y}$
+    $f(x,y)+f(x,y^{(n-1)})\geq f(x,y^{(n)})+f(x^{(n)},y^{(n-1)}),$
     referred to as Five-Point property, then the optimality of the limit is ensured.
-  • In [38], Grippo and Sciandrone prove convergence of the sequence to a stationary point of $f$, under the assumption of convexity of the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ and existence of the sequence limit.
-BA-type algorithms, introduced for the numerical computation of channel capacity curve [39] and RD curve [22], represent specific instances of alternating minimization algorithms [40, Chapter 9].
+  • In [40], Grippo and Sciandrone prove the convergence of the sequence to a stationary point of $f$, under the assumption of convexity of the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ and existence of the sequence limit.
+BA-type algorithms, introduced for the numerical computation of channel capacity [41] and RD function [24], are specific instances of alternating minimization algorithms [42, Chapter 9].
 In fact, in their classic formulation, both problems can be expressed as constrained minimization of a convex function on the sets of marginal distributions and transition matrices, where the properties of the sets (e.g., convexity) depend on the constraints for which the problem is formulated.
@@
 In this section, we present the derivation of our main theoretical results.
-We start by providing in the following lemma the parametric characterization of the solution of the RDPF problem, obtained by casting (1) as a double minimization problem.
+We start by providing the parametric characterization of the solution of the RDPF problem in the following lemma, obtained by casting (1) as a double minimization problem.
 Lemma 1.
@@
 Moreover, let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$ being the Lagrangian multipliers associated with constraints (2) and (3).
-Then (1) can be expressed as a double minimum as
-$\begin{split}R(D,P)=\min_{Q_{Y|X}\in\mathcal{Q}(\mathcal{X})}\min_{h\in\mathcal{P}(\mathcal{X})}D_{KL}(p_{X}Q_{Y|X}||p_{X}h)&+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}(D_{f}(p_{X}||q_{Y})-P)\end{split}$ (4)
-where $D=\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]$ and $P=D_{f}(p_{X}||q_{Y}^{*})$, with $Q_{Y|X}^{*}$ achieving the minimum.
-Furthermore, for fixed $Q_{Y|X}$, the right side of (4) is minimized by
+Then (1) can be expressed as a double minimum
+$\begin{split}R(D,P)=\min_{\begin{subarray}{c}Q_{Y|X}\in\mathcal{Q}(\mathcal{X})\\ h\in\mathcal{P}(\mathcal{X})\end{subarray}}&D_{KL}(p_{X}Q_{Y|X}||p_{X}h)\\ &+s_{D}\left(\mathbb{E}\left[d(X,Y)\right]-D\right)\\ &+s_{P}(D_{f}(p_{X}||q_{Y})-P)\end{split}$ (4)
+where $D=\mathbb{E}_{Q_{Y|X}^{*}}\left[d(X,Y)\right]$ and $P=D_{f}(p_{X}||h^{*})$, with $(Q_{Y|X}^{*},h^{*})$ being the pair achieving the minimum.
+Furthermore, for fixed $Q_{Y|X}$, the right-hand side of (4) is minimized by
 $h[Q_{Y|X}](y)=\sum_{x\in\mathcal{X}}p_{X}(x)Q_{Y|X}(y|x)$ (5)
-while, for fixed $h$, the right side of (4) is minimized by
+whereas, for fixed $h$, the right-hand side of (4) is minimized by
 $Q_{Y|X}[h](y|x)$ $=\frac{h(y)\cdot A[q_{Y}[h]](x,y,s)}{\sum_{i\in\mathcal{X}}h(i)\cdot A[q_{Y}[h]](x,i,s)}$ (6)
@@
 See Appendix A. ∎
-We remark that, although showing a close resemblance to the classical BA solution [41, Theorem 6.3.3], Lemma 1 differs from it in (6).
+We remark that, although showing a close resemblance to the classical BA solution [43, Theorem 6.3.3], Lemma 1 differs from it in (6).
 In particular, the perception constraint (3) induces the presence of an additional exponential term, i.e., $s_{P}g(\cdot,\cdot)$.
-Clearly, the classical BA implicit solution can be obtained as a special case of (6) by considering $s_{P}=0$, effectively removing the perceptual constraint.
+Note that the classical BA implicit solution can be obtained as a special case of (6) by considering $s_{P}=0$, effectively removing the perceptual constraint.
 The next corollary follows as a direct consequence of Lemma 1.
 Corollary 1.
-$R(D,P)$ in (4) can be reformulated as follows
-$\begin{split}R(D_{s},P_{s})=-s_{D}D_{s}-s_{P}P_{s}+\min_{h\in\mathcal{P}(\mathcal{X})}&\quad s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)\\ &-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[q_{Y}[h]](x,y,s)\right)\end{split}$ (9)
-where
-$D_{s}$ $=\sum_{(x,y)\in\mathcal{X}^{2}}\frac{p_{X}(x)h^{*}(y)A[h^{*}](x,y,s)}{\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)}d(x,y)\qquad P_{s}=D_{f}(p_{X}||h^{*})$
+Let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$. Then $R(D,P)$ in (4) can be reformulated as follows
+$R(D_{s},P_{s})=$ $-s_{D}D_{s}-s_{P}P_{s}$ (9)
+$+\min_{h\in\mathcal{P}(\mathcal{X})}$ $\quad s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)$
+$-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[q_{Y}[h]](x,y,s)\right)$
+where $P_{s}=D_{f}(p_{X}||h^{*})$ and
+$D_{s}$ $=\sum_{(x,y)\in\mathcal{X}^{2}}\frac{p_{X}(x)h^{*}(y)A[h^{*}](x,y,s)}{\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)}d(x,y)$
 with $h^{*}\in\mathcal{P}(\mathcal{X})$ achieving the minimum of (9).
 Proof:
-The proof follows by substitution of (6) in (4). ∎
+The proof follows by substitution of (6) into (4). ∎
+We note that in Corollary 1 and in subsequent analysis, the subscript notation $(D_{s},P_{s})$ is introduced to explicitly indicate the dependence of the constraint levels $(D,P)$ on the fixed Lagrangian multipliers $s=(s_{D},s_{P})$.
 The following lemma characterizes a necessary and sufficient condition to ensure that, for given Lagrangian multipliers $s=(s_{D},s_{P})$, a distribution $h^{*}\in\mathcal{P}(\mathcal{X})$ is the optimal solution of (9), i.e., $(h^{*},Q_{Y|X}[h^{*}])$ defines a point achieving the RDPF.
@@
 (Implicit Iterate) Due to the structure of the iterations in (6) and (8), an implicit dependency of $h^{(n+1)}$ on itself appears, i.e.,
-$\frac{h^{(n+1)}(y)}{h^{(n)}(y)}=\sum_{x\in\mathcal{X}}\frac{p_{X}(x)\exp\left\{-s_{D}d(x,y)-s_{P}g\left(p_{X},{h^{(n+1)}},y\right)\right\}}{\sum_{i\in\mathcal{X}}h^{(n)}(i)\exp\left\{-s_{D}d(x,i)-s_{P}g\left(p_{X},{h^{(n+1)}},i\right)\right\}}=c[h^{(n)},h^{(n+1)}](y)$ (11)
+$\frac{h^{(n+1)}(y)}{h^{(n)}(y)}$ $=\sum_{x\in\mathcal{X}}\frac{p_{X}(x)e^{-s_{D}d(x,y)-s_{P}g\left(p_{X},{h^{(n+1)}},y\right)}}{\sum_{i\in\mathcal{X}}h^{(n)}(i)e^{-s_{D}d(x,i)-s_{P}g\left(p_{X},{h^{(n+1)}},i\right)}}$
+$=c[h^{(n)},h^{(n+1)}](y)$ (11)
 showing that the updated term $h^{(n+1)}$ exists in both the left- and the right-hand side of the equation.
@@
 The implicit definition of $h^{(n+1)}$ in (11) suggests the application of numerical methods for its approximation.
-To this end, we introduce a variation to the OAM scheme, referred to as NAM scheme, where $h^{(n+1)}$ is approximated at each minimization step using Newton’s root-finding method [32].
+To this end, we introduce a variation of the OAM scheme, referred to as NAM scheme, where $h^{(n+1)}$ is approximated at each minimization step using Newton’s root-finding method [33].
 We first demonstrate that the iteration step for $h^{(n+1)}$, i.e., (11), can be cast as a root finding problem.
@@
 The proof follows from the evaluation of (12) in $h^{(n+1)}$ and the substitution of (11) therein. ∎
-The application of Newton’s method requires the existence and the invertibility of the Jacobian matrix $J_{T}$ of the functional $T$ [32, Section 10.2].
+The application of Newton’s method requires the existence and the invertibility of the Jacobian matrix $J_{T}$ of the functional $T$ [33, Section 10.2].
 In our case, ensuring the existence of $J_{T}$ requires a more restrictive continuity assumption on the divergence measure, i.e., $D_{f}(\cdot||\cdot)$ needs to be twice differentiable in its second argument.
-Although this limitation reduces the generality of the NAM scheme, we note that the majority of the commonly used divergences, i.e., see Section II-B, respect this assumption.
+Although this limitation reduces the generality of the NAM scheme, we note that most commonly used divergences (see Section II-B) satisfy this assumption.
 Under this restriction, the invertibility of $J_{T}$ is shown in the following lemma.
@@
 where
-$M[h^{(n)},u]$
+$M[h^{(n)},u]=$ (14)
+$~\left[h^{(n)}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[u](x,i,s)\cdot A[u](x,j,s)}{\left(\sum_{k\in\mathcal{X}}h^{(n)}(k)A[u](x,k,s)\right)^{2}}\right]$
 $\Gamma[h^{(n)},u]$ $=s_{P}\operatorname{diag}\left[h^{(n)}(i)\cdot\frac{\partial^{2}D_{f}(p_{X}||v)}{\partial v(i)^{2}}\bigg{|}_{u}\right]_{i\in\mathcal{X}}$ (15)
@@
 Furthermore, let the sequence $\{u^{(k)}\}_{k=1,2,\ldots}$ for some initial point $u^{(0)}\in\mathbb{R}^{|\mathcal{X}|}$ be defined as
-$u^{(k+1)}\triangleq u^{(k)}-\left(J_{T}[h^{(n)},u^{k}]\right)^{-1}\cdot T[h^{(n)},u^{(k)}].$
+$u^{(k+1)}\triangleq u^{(k)}-\left(J_{T}[h^{(n)},u^{(k)}]\right)^{-1}\cdot T[h^{(n)},u^{(k)}].$
 Then, $\lim_{k\to\infty}u^{(k)}=h^{(n+1)}$.
 Proof:
-The proof follows by direct application of Newton’s root-finding method [32, Section 10.2], since Lemma 3 proves that the set of solutions $h^{(n+1)}$ and the set of the roots of $T(\cdot)$ coincide, while Lemma 4 proves that $T(\cdot)$ satisfies the assumption for the convergence of the method.
+The proof follows by direct application of Newton’s root-finding method [33, Section 10.2], since Lemma 3 proves that the set of solutions $h^{(n+1)}$ and the set of the roots of $T(\cdot)$ coincide, while Lemma 4 proves that $T(\cdot)$ satisfies the assumption for the convergence of the method.
 ∎
@@
 where $A[\cdot]$ is defined in (7) and $v[\cdot]:\mathcal{P}(\mathcal{X})\to\mathcal{P}(\mathcal{X})$ is any functional defining a probability distribution.
-Let $\hat{h}^{(0)}$ be any probability vector with nonzero components and let $\hat{Q}_{Y|X}^{(n+1)}\equiv\hat{Q}_{Y|X}[\hat{h}^{(n)}]$, $\hat{h}^{(n+1)}\equiv\hat{q}_{Y}[\hat{h}^{(n)}]$, and $v^{(n)}=v[h^{(n)}]$.
+Let $\hat{h}^{(0)}$ be any probability vector with nonzero components and let $\hat{Q}_{Y|X}^{(n+1)}\equiv\hat{Q}_{Y|X}[\hat{h}^{(n)}]$, $\hat{h}^{(n+1)}\equiv\hat{q}_{Y}[\hat{h}^{(n)}]$, and $v^{(n)}=v[\hat{h}^{(n)}]$.
 Then, as $n\xrightarrow{}\infty$, we obtain
 $D(\hat{Q}_{Y|X}^{(n)})\xrightarrow{}D_{s},~P(\hat{Q}_{Y|X}^{(n)})$ $\xrightarrow{}P_{s},~I(p_{X},\hat{Q}_{Y|X}^{(n)})\xrightarrow{}R(D_{s},P_{s})$
-if $\lim_{n\rightarrow\infty}||h^{(n+1)}-v^{(n)}||=0$ with at least linear rate of convergence.
+[…]
 Proof:
@@
 Theorem 4 enables the implementation of the alternating minimization scheme by introducing an auxiliary variable $v[h^{(n)}]$, which approximates the correct iteration $h^{(n+1)}$ while still being a function of only the current iteration of $h^{(n)}$.
-Nevertheless, depending on $v$, this approximation may incur restrictions on the domain of the Lagrangian multiplier $s_{P}$ that affect convergence guarantees, as we will later discuss in Section IV.
+Nevertheless, depending on $v$, this approximation may incur restrictions on the domain of the Lagrangian multiplier $s_{P}$ that affect convergence guarantees, as discussed later in Section IV.
 We conclude this section with the following technical remark, highlighting the differences between the NAM and RAM schemes.
@@
 (NAM vs.
-RAM) The main advantage of NAM is the convergence for any value of the Lagrangian multipliers $(s_{D},s_{P})$, without needing any additional condition.
+RAM) The main advantage of NAM is the convergence for any value of the Lagrangian multipliers $(s_{D},s_{P})$, without the need for any additional condition.
 However, the introduction of Newton’s method requires the differentiability of the perceptual metric and the introduction of additional complexity at each iteration.
-RAM, on the other hand, lifts the differentiability requirement and avoids the additional computational cost at each iteration, but at the expense of a potentially smaller set of $(s_{D},s_{P})$ for which the algorithm achieves the optimal solution, potentially prohibiting the computation of the complete RDP curve.
+RAM, on the other hand, removes the differentiability requirement and avoids the additional computational cost at each iteration, but at the expense of a potentially smaller set of $(s_{D},s_{P})$ for which the algorithm achieves the optimal solution, which may preclude the computation of the complete RDP curve.
 IV Algorithmic Implementation and Convergence Analysis
-This section is about the algorithmic implementation of the alternating minimization schemes derived in Section III and the characterization of their convergence rate.
+This section addresses the algorithmic implementation of the alternating minimization schemes derived in Section III and the characterization of their convergence rate.
 We start by presenting the implementation of the NAM and RAM schemes, respectively, in Algorithm 1 and 2. Subsequently, we discuss the derivation of stopping conditions suitable for both algorithms.
 Algorithm 1 Newton-based Alternating Minimization (NAM)
-1:source distribution $p_{X}$; Lagrangian parameters $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\geq 0$; error tolerances $\epsilon>0$, $\delta>0$; distortion measure $d(x,y)$; initial assignment $h^{(0)}$.
-2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
-3:while $\omega\geq\epsilon$ do
-4:     $h^{(n+1)}\leftarrow$ Newton Approximation( $p_{X}$, $h^{(n)}$, $s$, $\delta$)
-5:     $c^{(n)}\leftarrow c\left[h^{(n)},h^{(n+1)}\right]$
-6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}h^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
-7:     $n\leftarrow n+1$
-8:end while
+1:source distribution $p_{X}$; Lagrangian parameters $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\geq 0$; error tolerances $\epsilon>0$, $\delta>0$; distortion measure $d(\cdot,\cdot)$; initial assignment $h^{(0)}$.
+2:
+[…]
+4:while $\omega\geq\epsilon$ do
+5:  $h^{(n+1)}\leftarrow$ Newton Approx.( $p_{X}$, $h^{(n)}$, $s$, $\delta$)
+6:  $c^{(n)}\leftarrow c\left[h^{(n)},h^{(n+1)}\right]$
+7:  $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}h^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
+8:  $n\leftarrow n+1$
+9:end while
+10:
 […]
@@
 1:source distribution $p_{X}$; Lagrangian multipliers $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\in[0,s_{P,\max}]$; error tolerance $\epsilon>0$; divergence measure $D_{f}(\cdot||\cdot)$; distortion measure $d(\cdot,\cdot)$; initial assignment $\hat{h}^{(0)}$.
-2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
-3:while $\omega\geq\epsilon$ do
-4:     $c^{(n)}\leftarrow c[\hat{h}^{(n)},v^{(n)}]$
-5:     $\hat{h}^{(n+1)}\leftarrow\hat{h}^{(n)}\cdot c^{(n)}$
-6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}\hat{h}^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
-7:     $n\leftarrow n+1$
-8:end while
+2:
+[…]
+4:while $\omega\geq\epsilon$ do
+5:  $c^{(n)}\leftarrow c[\hat{h}^{(n)},v^{(n)}]$
+6:  $\hat{h}^{(n+1)}\leftarrow\hat{h}^{(n)}\cdot c^{(n)}$
+7:  $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}\hat{h}^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
+8:  $n\leftarrow n+1$
+9:end while
+10:
 […]
 Stopping Criterion
-We first focus on the derivation of a stopping criterion for the RAM case, as the NAM case can be obtained by fixing the auxiliary variable $v[h^{(n)}]=q_{Y}[h^{(n)}]$ in Theorem 4, i.e., recovering the original OAM iterates.
-For this purpose, we need the following theorem in which we derive bounds on the RDPF.
+We first derive a stopping criterion for the RAM case, since the NAM case can be obtained by fixing the auxiliary variable $v[h^{(n)}]=q_{Y}[h^{(n)}]$ in Theorem 4, i.e., recovering the original OAM iterates.
+For this purpose, we need the following theorem which establishes bounds on the RDPF.
 Theorem 5.
-(Bounds on RDPF) Let $\hat{Q}_{Y|X}$ and $\hat{h}$ be defined as in Theorem 4, $c()$ be as defined as in Lemma 2, and $c_{\max}=\max_{y\in\mathcal{X}}c[u,\hat{q}_{Y}[u]](y)$.
-Then, at the point $D=\mathbb{E}_{\hat{Q}_{Y|X}}\left[d(x,y)\right]$, and $P=D_{f}(p_{X}||\hat{q}_{Y})$, the following bounds hold
-$R(D,P)$ $\geq R_{L}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]-\log(c_{\max})$ (17)
-$\begin{split}R(D,P)&\leq R_{U}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\sum_{y\in\mathcal{X}}u(y)c[u,\hat{q}_{Y}[u]](y)\log(c[u,\hat{q}_{Y}[u]](y))\end{split}$ (18)
-where $\hat{W}{}[\hat{h}]$ is given by
-$\begin{split}\hat{W}[u]&=-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}u(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{u(y)}\partial f\left(\frac{p_{X}(y)}{u(y)}\right)\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u]\left(f\left(\frac{p_{X}(y)}{q_{Y}[u](y)}\right)-f\left(\frac{p_{X}(y)}{u(y)}\right)\right)\Bigg{]}.\end{split}$ (19)
+[…]
+$\begin{split}R(D,P)&\geq R_{L}[\hat{h}](D,P)\\ &=-s_{D}D-s_{P}P+\hat{W}[\hat{h}]-\log(c_{\max})\end{split}$ (18)
+$\begin{split}R(D,P)&\leq R_{U}[\hat{h}](D,P)\\ &=-s_{D}D-s_{P}P+\hat{W}[\hat{h}]\\ &~~~~-\sum_{y\in\mathcal{X}}\hat{h}(y)c[\hat{h},\hat{q}_{Y}[\hat{h}]](y)\log(c[\hat{h},\hat{q}_{Y}[\hat{h}]](y))\end{split}$ (19)
+where $\hat{W}{}$$\cdot$] is given by
+$\begin{split}\hat{W}[u]&=-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}u(y)A[v[u]](x,y,s)\right)\\ &\quad+s_{P}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u](y)\frac{p_{X}(y)}{v[u](y)}\partial f\left(\frac{p_{X}(y)}{v[u](y)}\right)\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u]\left(f\left(\frac{p_{X}(y)}{\hat{q}_{Y}[u](y)}\right)-f\left(\frac{p_{X}(y)}{v[u](y)}\right)\right)\Bigg{]}.\end{split}$ (20)
 Proof:
@@
 OAM Convergence Rate
-We remark that the iteration structure in Theorem 2, i.e., $h^{(n+1)}$ = $q_{Y}[h^{(n)}]$ given the current iteration $n$, can be described as an implicit vector function $S:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ with $S[h](i)=h(i)\cdot c[h,S[h]](i)$, such that $h^{(n+1)}=S[h^{(n)}]$.
-The results of Lemma 2 characterize a distribution $h^{*}$ achieving the RDPF as a fixed point of $S(h)$, i.e.
-$h^{*}=S[h^{*}]$, since $c[h^{*},S[h^{*}]](i)=1,~{}i=1\ldots,|\mathcal{X}|$.
+We note that the iteration structure in Theorem 2, i.e., $h^{(n+1)}$ = $q_{Y}[h^{(n)}]$ given the current iteration $n$, can be represented as an implicit vector function $S:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ with $S[h](i)=h(i)\cdot c[h,S[h]](i)$, such that $h^{(n+1)}=S[h^{(n)}]$.
+The results of Lemma 2 characterize a distribution $h^{*}$ that achieves the RDPF as a fixed point of $S(h)$, i.e., $h^{*}=S[h^{*}]$, since $c[h^{*},S[h^{*}]](i)=1,~i=1\ldots,|\mathcal{X}|$.
 Under these observations, we can analyze the convergence rate of the OAM scheme following similar steps as in [44].
-The first order Taylor expansion of $S[h]$ around a fixed point $h^{*}$ is defined as
-$S[h]=S[h^{*}]+J[h^{*}]\cdot(h-h^{*})+o(||h-h^{*}||)$
+The first-order Taylor expansion of $S[h]$ around a fixed point $h^{*}$ is defined as
+[…]
 where $J[h]$ is the Jacobian matrix of $S[h]$ with entries $J[h](i,j)\triangleq\frac{\partial S[h](i)}{\partial h(j)},(i,j)\in\mathcal{X}^{2}$.
@@
 Theorem 6.
-(Jacobian form) The Jacobian $J(h)$ computed at the fixed point $h^{*}$ is given as
+(Jacobian form) The Jacobian $J(h)$ evaluated at the fixed point $h^{*}$ is given as
 $J[h^{*}]=\left(I-M^{*}\right)\left(I-\Gamma^{*}J[h^{*}]\right)$ (21)
-where $M^{*}=M[h^{*},h^{*}]$ and $\Gamma^{*}=\Gamma[h^{*},h^{*}]$ as defined in (LABEL:matrix:_Newton_M) and (14), respectively.
+where $M^{*}=M[h^{*},h^{*}]$ and $\Gamma^{*}=\Gamma[h^{*},h^{*}]$ as defined in (14) and (15), respectively.
 Proof:
@@
 Let $\{\lambda_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the set of eigenvalues of $M^{*}=M[h^{*},h^{*}]$.
-Given a distortion function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ that induces a full-rank matrix $D=[e^{-s_{D}d(i,j)}]_{(i,j)\in\mathcal{X}^{2}}$, then $\lambda_{i}>0$,$\forall i\in[1:|\mathcal{X}|]$, i.e., $M^{*}$ has only positive eigenvalues.
+Given a distortion function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ that induces a full-rank matrix $D=[e^{-s_{D}d(i,j)}]_{(i,j)\in\mathcal{X}^{2}}$, then $\lambda_{i}>0$$\forall i\in[1:|\mathcal{X}|]$, i.e., $M^{*}$ has only positive eigenvalues.
 Proof:
@@
 (Convergence rate of Theorem 2) Let $\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the eigenvalues of $J[h^{*}]$. Then,
-$1\displaystyle 0\leq\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}<$ 1 .
-Moreover, let $\gamma\in[\theta_{\max},1)$. Then, there exists $\delta>0$ and $K>0$ such that if $h^{(0)}\in\{h:||h-h^{*}||\leq\delta\}$, we obtain
-$||h^{(n)}-h^{*}||<K\cdot||h^{(0)}-h^{*}||\cdot\gamma^{n}$ (21)
+$0\leq\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}<1.$
+[…]
+[…]
 i.e., the iterations converge exponentially.
@@
 NAM Convergence Rate
-The convergence rate of the NAM scheme is immediate from the OAM scheme analysis, given the close relation between the two schemes.
-Since the only difference lies in the introduction of Newton’s root-finding method for the estimation of the optimal iteration step, the NAM scheme enjoys the same convergence rate in terms of the number of iterations as the OAM scheme, i.e., an exponential convergence $\mathcal{O}(e^{-n})$ under the assumptions of Lemma 5.
+The convergence rate of the NAM scheme follows directly from the OAM scheme analysis, given the close relation between the two schemes.
+Since the only difference lies in the introduction of Newton’s root-finding method for the estimation of the optimal iteration step, the NAM scheme exhibits the same convergence rate in terms of the number of iterations as the OAM scheme, i.e., an exponential convergence $\mathcal{O}(e^{-n})$ under the assumptions of Lemma 5.
 However, the added complexity from the application of Newton’s method at each iteration increases the overall iteration complexity, due to the at least linear convergence rate $\mathcal{O}(\frac{1}{m})$ of the root approximant.
@@
 $\hat{J}[h^{*}]=(I-M[h^{*},h^{*}])(I-\Gamma[h^{*},h^{*}])$
-where $M$ and $\Gamma$ are given by (LABEL:matrix:_Newton_M) and (14), respectively.
+where $M$ and $\Gamma$ are given by (14) and (15), respectively.
 Unlike Theorem 2, where the structure of (21) bounds its own eigenvalues, in this case, we need to bound the Lagrangian multiplier $s_{P}$, hence matrix $\Gamma$, to guarantee exponential convergence of the algorithm.
@@
 Define the set $I_{s_{P}}^{\epsilon}=[0,s_{P,\max}-\epsilon]$ for $0<\epsilon<s_{P,\max}$.
-Then, there exists an $\epsilon^{\prime}$ such that if $s_{P}\in I_{s_{P}}^{\epsilon^{\prime}}$ then $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1.
+Then, there exists an $\epsilon^{\prime}$ such that if $s_{P}\in I_{s_{P}}^{\epsilon^{\prime}}$ then $0\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<1$.
 Proof:
@@
 V Numerical Results
-In this section, we validate our theoretical findings with simulations. In particular, we consider the computation of the RDPF under the NAM and RAM schemes, respectively using Algorithms 1 and 2.
+In this section, we validate our theoretical findings through simulations. In particular, we consider the computation of the RDPF under the NAM and RAM schemes, using Algorithms 1 and 2, respectively.
 V-A RDPF Computation - NAM scheme
 Example 1.
-Suppose that ${\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(\cdot||\cdot)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{4}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$.
+Suppose that ${\mathcal{X}}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with the perception constraint chosen as one of the following: (a) $D_{f}(\cdot||\cdot)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$.
 In Fig.
-1, we display the $R(D,P)$ estimates resulting from Algorithm 1 for each divergence metric.
+1, we present the $R(D,P)$ estimates obtained using Algorithm 1 for each divergence metric.
 (a) $D_{JS}(\cdot||\cdot)$
@@
 (d) $D_{\alpha=-1}$
-(e) $D_{\alpha=\frac{1}{4}}$
-(f) $D_{\alpha=\frac{1}{2}}$
-Figure 1: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
-We observe that all the computed RDPFs share similarities in the structure of the operating regions on the $(D,P)$ plane. Taking as a reference the plot of Fig. 1(a), we distinguish three cases:
-  • Case I, where the perception level $P$ is not met with equality. In this region, the RDPF is equal to the RDF.
+(e) $D_{\alpha=\frac{1}{2}}$
+Figure 1: $R(D,P)$ for a Bernoulli source under Hamming distortion and various perception constraints.
+We observe that all the computed RDPFs share similarities in the structure of the operating regions on the $(D,P)$ plane. Referring to the plot of Fig. 1a, we distinguish three cases:
+  • Case I, where the perception constraint $P$ is not met with equality. In this region, the RDPF is equal to the RDF.
   • Case II, where, due to the distortion level $D$ not met with equality, the RDPF function is identically zero.
@@
 The results show that Algorithm 1 can completely cover the Case III region.
-The boundary between Cases I-III (black line) representing the RDF is obtained by setting the Lagrangian multiplies $s_{P}=0$.
+The boundary between Cases I-III (black line) representing the RDF is obtained by setting the Lagrangian multipliers $s_{P}=0$.
 Furthermore, the Case I region can be obtained by extension of the relative boundaries, since all the $(D,P)$ points share the solution given by the RD problem.
@@
 Example 2.
-Suppose that ${\cal X}=\hat{\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(p_{X}||h)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=TV(\cdot||\cdot)$.
+Suppose that ${\mathcal{X}}=\hat{\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(p_{X}||h)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=TV(\cdot||\cdot)$.
 In Fig.
-2, we display the $R(D,P)$ estimates resulting from Algorithm 2 for each divergence metric.
+2, we present the $R(D,P)$ estimates obtained using Algorithm 2 for each divergence metric.
 (a) $D_{JS}(\cdot||\cdot)$
@@
 (d) $D_{\alpha=-1}(\cdot||\cdot)$
-(e) $D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$
-(f) $TV(\cdot||\cdot)$
-Figure 2: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
-Remarks similar to those in Example 1 can be made regarding the operating regions of the computed RDPFs.
+(e) $TV(\cdot||\cdot)$
+Figure 2: $R(D,P)$ for a Bernoulli source under Hamming distortion and various perception constraints.
+Similar remarks to those in Example 1 apply to the operating regions of the computed RDPFs.
 In addition, in Fig.
-2(f) we compare the theoretical results of [3, Equation 6] with the numerical results obtained using Algorithm 2.
+2e we compare the theoretical results of [3, Equation 6] with the numerical results obtained using Algorithm 2.
 This case is of particular interest due to the non-differentiability of the $TV$ distance.
 We observe that Algorithm 2 achieves exactly the theoretical solution of [3, Equation 6] as long as $D\leq{D}_{\max}=0.15$.
-We attribute this behavior to the limitations in the domain of the Lagrangian multiplier $s_{P}$, as discussed in Theorem 8.
+We attribute this behavior to limitations in the domain of the Lagrangian multiplier $s_{P}$, as discussed in Theorem 8.
 We address this issue in detail in the following section.
 V-C On the convergence under TV perception constraint
-As observed in Example 2, in the case of TV distance the RAM algorithm is not able to converge to proper solutions in the region $\Omega=\{(D,P)\in\mathbb{R}_{+}^{2}:D\geq p=0.15\}$.
-We argue that the problem stems from the values of the Lagrangian multipliers $s=(s_{D},s_{P})$ associated with $\Omega$ requires values of $s_{P}$ that do not guarantee the convergence of the algorithm, as reported in Theorem 8.
+As observed in Example 2, in the case of TV distance, the RAM algorithm is not able to converge to proper solutions in the region $\Omega=\{(D,P)\in\mathbb{R}_{+}^{2}:D\geq p=0.15\}$.
+We argue that the problem stems from the values of the Lagrangian multipliers $s=(s_{D},s_{P})$ associated with $\Omega$ require values of $s_{P}$ that do not guarantee the convergence of the algorithm, as reported in Theorem 8.
 To solve the issue, we propose an approximation of the $TV(\cdot||\cdot)$ distance through a sequence of $f$-divergences $\{D_{f_{n}}\}_{n\in\mathbb{N}}$ such that ${D_{f_{n}}}\to TV$ for $n\to\infty$.
@@
 For a divergence metric $D_{f}(\cdot||\cdot)$, let the set $\mathcal{L}_{D_{f}}(D,P)$ be defined as
-$\mathcal{L}_{D_{f}}(D,P)\triangleq\{Q_{Y|X}\in\mathcal{Q}:\mathbb{E}_{Q_{Y|X}}\left[d(x,y)\right]\leq D\land D_{f}(p_{X}||q_{Y})\leq P\}.$
+$\mathcal{L}_{D_{f}}(D,P)\triangleq\{Q_{Y|X}:\mathbb{E}_{Q_{Y|X}}\left[d(X,Y)\right]\leq D,D_{f}(p_{X}||q_{Y})\leq P\}.$
 Given $D_{f}$, $D_{g}$ divergence metrics with $D_{f}(p||q)\leq D_{g}(p||q),\forall p,q\in\mathcal{P}$, then $\mathcal{L}_{D_{g}}(D,P)\subseteq\mathcal{L}_{D_{f}}(D,P)$.
 Moreover, for the associated RDPF problems
-$R_{D_{f}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{f}}(D,P)}I(p_{X},Q_{Y|X})\quad\text{and}\quad R_{D_{g}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{g}}(D,P)}I(p_{X},Q_{Y|X})$
+$R_{D_{f}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{f}}(D,P)}I(p_{X},Q_{Y|X})$
+$R_{D_{g}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{g}}(D,P)}I(p_{X},Q_{Y|X})$
 the inequality $R_{D_{g}}(D,P)\geq R_{D_{f}}(D,P)$ holds.
@@
 See Appendix L. ∎
-The results from Lemmas 7 and 8 guarantee that, $\forall n\in\mathbb{N}$, the RDP problem defined for the perception metric $D_{f_{n}}(\cdot||\cdot)$ act as lower-bound for the RDP problem defined with $TV(\cdot||\cdot)$ perception metric.
-Furthermore, since $D_{f_{n}}(\cdot||\cdot)$ is a smooth function, we can apply NAM scheme to compute the associate RDPF.
+The results from Lemmas 7 and 8 guarantee that, $\forall n\in\mathbb{N}$, the RDP problem defined for the perception metric $D_{f_{n}}(\cdot||\cdot)$ acts as a lower bound for the RDP problem defined with $TV(\cdot||\cdot)$ perception metric.
+Furthermore, since $D_{f_{n}}(\cdot||\cdot)$ is a smooth function, we can apply the NAM scheme to compute the associated RDPF.
 Fig.
-3 shows the RDPFs for a source $X\sim Ber(0.15)$ under Hamming distortion and perception measure $D_{f_{n}}$, for $n=\{1,10,100\}$.
-As expected, for increasing $n$, the $D_{f_{n}}$ perception metric becomes a better approximation of the $TV(\cdot||\cdot)$ perception.
+3 shows the RDPFs for a source $X\sim Ber(0.15)$ under Hamming distortion and perception measure $D_{f_{n}}$, for $n\in\{1,10,100\}$.
+As expected, for increasing $n$, $D_{f_{n}}$ provides a progressively better approximation of the $TV(\cdot||\cdot)$ metric.
 (a) $TV(\cdot||\cdot)$
@@
 (d) $n=100$
-Figure 3: Comparison between the RDPF under $TV$ perception, computed with RAM scheme, and the RDPF under $D_{f_{n}}$, computed with NAM scheme, for $n$={1,10, 100}.
+Figure 3: Comparison between the RDPF under TVTV perception, computed with the RAM scheme, and the RDPF under $D_{f_{n}}$, computed with the NAM scheme, for $n\in\{1,10,100\}$.
 VI Conclusion and Future Directions
-In this paper, we studied the computation of the rate-distortion-perception function for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences.
-We showed that the introduction of the perception constraint induces a non-trivial generalization of the classical optimality conditions in terms of reconstruction distribution and transition matrix $(h^{*},Q_{Y|X}^{*})$, resulting in the OAM scheme.
-The presence of implicit equations in the formulation of the OAM scheme, implying the impossibility of direct algorithmic implementation, led to the formulation of alternative minimization schemes, NAM and RAM, each with its applicability conditions.
+In this paper, we have studied the computation of the rate-distortion-perception function for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint from the family of $f$-divergences.
+We showed that the introduction of the perception constraint induces a non-trivial generalization of the classical optimality conditions for the reconstruction distribution and the transition matrix $(h^{*},Q_{Y|X}^{*})$, resulting in the OAM scheme.
+Since the OAM scheme involves implicit equations that preclude direct algorithmic implementation, we introduced alternative minimization schemes, NAM and RAM, each with its own applicability conditions.
 Furthermore, we analyzed the asymptotic complexity of the proposed schemes and provided numerical results to validate our theoretical findings.
-As an ongoing research, we plan to extend the derived optimality conditions to general measure spaces, as well as to study the behavior of the proposed algorithm when applied to continuous sources.
+As ongoing research, we plan to extend the derived optimality conditions to general measure spaces, as well as to study the behavior of the proposed algorithms when applied to continuous sources.
 Appendix A Proof of Lemma 1
-The double minimization follows immediately from properties of the KL divergence, reported in [41, Theorem 5.2.6], which allows to rewrite (1) as (4).
+The double minimization follows immediately from properties of the KL divergence, reported in [43, Theorem 5.2.6], which allows to rewrite (1) as (4).
 Moreover, the characterization (5) of the minimizer $h$ follows from the same properties.
-On the other hand, the minimizing $Q_{Y|X}$ can be characterized by applying the Karush-Kuhn-Tucker (KKT) conditions [44] on the dual formulation of (1).
-The Lagrangian associated with the primal problem has form:
-$\begin{split}L(Q_{Y|X},s,\lambda,\mu)&=D_{KL}\left(p_{X}Q_{Y|X}||p_{X}h\right)+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}\left(D_{f}(p_{X}||q_{Y})-P\right)\\ &\quad+\sum_{x\in\mathcal{X}}\lambda_{x}\left(1-\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)\right)-\sum_{(x,y)\in\mathcal{X}^{2}}\mu_{x,y}Q_{Y|X}(y|x)\end{split}$
+On the other hand, we remark that (1) is a convex program in the variable $Q_{Y|X}$ for a given $p_{X}$, and respects Slater’s condition, since it is easy to show that $Q_{Y|X}(y|x)=\delta_{y=x}$ is an interior point of the constraint set, as highlighted in Remark 2.
+Therefore, the minimizer $Q_{Y|X}^{*}$ can be characterized by applying the Karush-Kuhn-Tucker (KKT) conditions [46] on the dual formulation of (1).
+The Lagrangian associated with the primal problem has the form:
+$L(Q_{Y|X},s,\lambda,\mu)$
+$=D_{KL}\left(p_{X}Q_{Y|X}||p_{X}h\right)+s_{D}\left(\mathbb{E}\left[d(X,Y)\right]-D\right)$
+$\quad+s_{P}\left(D_{f}(p_{X}||q_{Y})-P\right)+\sum_{x\in\mathcal{X}}\lambda_{x}\left(1-\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)\right)$
+$\quad-\sum_{(x,y)\in\mathcal{X}^{2}}\mu_{x,y}Q_{Y|X}(y|x)$
 where the last two sets of constraints refer to the positivity and normalization constraints on $Q_{Y|X}$.
-Differentiating $L(Q_{Y|X},s,\lambda,\mu)$ with respect to the primal variables $Q_{Y|X}(y|x),\forall(x,y)\in\mathcal{X}^{2}$, we obtain:
-$\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}+s_{D}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}+s_{P}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}+\lambda_{x}-\mu_{x,y}\end{split}$
+Differentiating $L(Q_{Y|X},s,\lambda,\mu)$ with respect to the primal variables $Q_{Y|X}(y|x),\forall(x,y)\in\mathcal{X}^{2}$, we obtain
+$\begin{split}&\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}\\ &~~~~~~~~~=\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}+s_{D}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}\\ &~~~~~~~~~~~~~+s_{P}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}+\lambda_{x}-\mu_{x,y}\end{split}$
 where:
@@
 Enforcing stationarity and complementary slackness
-$\begin{cases}\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=0\\ \mu(x,y)Q_{Y|X}(y|x)=0\\ \end{split}\end{cases}$ (22)
+[…]
 we solve for $Q_{Y|X}(y|x)$ while choosing $\lambda(x)$ such that $\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)=1$, $\forall x\in\mathcal{X}$, resulting in (6) and (7). This completes the proof.
@@
 Using the results of Corollary 1, we can apply KKT conditions to (9). Thus, a minimum for $h$ must satisfy:
-$\begin{split}\frac{\partial}{h(y)}\Bigg{[}-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[h](x,y,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)+\lambda\sum_{y\in\mathcal{X}}h(y)\Bigg{]}\geq 0\end{split}$
+$\begin{split}\frac{\partial}{h(y)}&\Bigg{[}-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[h](x,y,s)\right)\\ &+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)+\lambda\sum_{y\in\mathcal{X}}h(y)\Bigg{]}\geq 0\end{split}$
 which reduces to
@@
 Let $\textbf{h}=(h^{(0)},h^{(1)},\ldots)$, and $\textbf{Q}=(Q_{Y|X}^{(1)},Q_{Y|X}^{(2)},\ldots)$ be the sequences of probability vectors and transition matrices obtained by the chain of alternating minimization $h^{(n)}\xrightarrow{}Q_{Y|X}^{(n+1)}\xrightarrow{}h^{(n+1)}$.
-Let $A^{(n)}=A[h^{(n)}]$ and define the functionals $V[h^{(n)}]$ and $W[h^{(n)}]$ as
-$\begin{split}&V[h^{(n)}]=D_{KL}\left(p_{X}\cdot Q_{Y|X}^{(n+1)}||p_{X}\cdot h^{(n)}\right)+s_{D}\mathbf{E}_{Q_{Y|X}^{(n+1)}p_{X}}[d(x,y)]+s_{P}D_{f}(p_{X}||h^{(n+1)})\end{split}$ (23)
-$\begin{split}&W[h^{(n)}]=s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h^{(n)}(y)A^{(n)}(x,,s)\right).\end{split}$ (24)
+Let $A^{(n)}=A[h^{(n+1)}]$ and define the functionals $V[h^{(n)}]$ and $W[h^{(n)}]$ as
+$V[h^{(n)}]=$ $D_{KL}\left(p_{X}\cdot Q_{Y|X}^{(n+1)}||p_{X}\cdot h^{(n)}\right)$ (24)
+$+s_{D}\mathbb{E}_{p_{X}Q_{Y|X}^{(n+1)}}\left[d(X,Y)\right]+s_{P}D_{f}(p_{X}||h^{(n+1)})$
+$W[h^{(n)}]$ $=s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$ (25)
+$~~~~-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h^{(n)}(y)A^{(n)}(x,,s)\right)$
 Using Theorem 1 and Corollary 1, we can show that $V[\textbf{q}]$ is non-increasing on the sequence q by first fixing $h^{(n)}$ and minimizing over Q (resulting in $Q_{Y|X}^{(n+1)}$) and then fixing $Q_{Y|X}^{(n+1)}$ and minimizing over q (resulting in $h^{(n+1)}$).
@@
 Let now $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and let
-$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)\\ &\quad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{\left(n\right)}(x,i,s)\right)\end{split}$ (25)
+$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)\\ &=\quad\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)\\ &\qquad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\qquad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{\left(n\right)}(x,i,s)\right)\end{split}$ (26)
 where
-$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)$ $=-s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]-s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
-$\quad+s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right).$
+$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)$
+$\quad=-s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(X,Y)\right]-s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
+$\qquad+s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right).$
 We can introduce an upper bound to (26) by noticing that
-$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}&p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}\cdot h^{*}}{h^{(n)}\cdot Q_{Y|X}^{*}}\right)\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{n}\log\left(\frac{h^{*}}{h^{(n)}}\right)-1=0.\end{split}$ (26)
-As a result, we obtain
-$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)\leq$
-$\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)+s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]+s_{P}D_{f}(p_{X}||h^{*})$
-$\quad+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)\right]$
-$\quad-\Biggl{[}-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{(n)}(x,i,s)\right)+s_{P}\sum_{y\in\mathcal{X}}h^{n+1}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
-$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}$
-$\leq V[h^{*}]-W[h^{(n)}]+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}(p_{X}||h^{*})\right]$
-$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}.$
-Notice that for any iteration $n\geq 0$, we have $W[h^{(n)}]\geq V[h^{(n+1)}]\geq V[h^{*}]$ and, subsequently, the following inequalities
-$0\leq W[h^{(n)}]-V[h^{*}]\leq G[h^{(n)}]$ (27)
-where
-$G[h^{(n)}]$ $=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)+s_{P}\left\{\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)\right\}$
-$\quad+s_{P}\Biggl{\{}\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]\Biggr{\}}.$
-Since $s_{P}\geq 0$ and due to the fact that
-$\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)+\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]$
-$=D_{f}(p_{X}||h^{(n+1)})-D_{f}(p_{X}||h^{(*)})-\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
-$\quad+\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)\frac{p_{X}(i)}{h^{(n+1)}}\partial f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
-$=D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||h^{(*)}\right)-\sum_{i\in\mathcal{X}}\left(h^{(n+1}(i)-h^{*}(i)\right)\frac{\partial D_{f}(p_{X}||u)}{\partial u(i)}\Bigg{|}_{h^{(n+1)}}$
-$=-\left(D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*})\right)\stackrel{{\scriptstyle(a)}}{{\leq}}0$
-where $T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*}))$ is the first order Taylor expansion of $D_{f}(p_{X}||\cdot)$, centered in $h^{(n+1)}$ and evaluated in $h^{(*)}$, and $(a)$ is verified since $D_{f}(p_{X}||\cdot)$ is a convex function in its second argument.
-Therefore $\forall n\in\mathbf{N}$, we obtain
-$W[h^{(n)}]-V[h^{*}]\leq$ $\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right).$
-Summing over $N$ terms, we obtain
-$\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ $\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)$
-$=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(N+1)}}{Q_{Y|X}^{(1)}}\right)$
-$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{Q_{Y|X}^{(1)}}\right)$
-where $(b)$ follows using the logarithm inequality Since at any iteration $n$, $W[h^{(n)}]-V[h^{*}]\geq 0$ and for all integers $N>0$ the partial sum is upper-bounded by a constant $L(h^{*},h^{(0)})$ dependent only on the initial probability assignment $h^{(0)}$, we have that $\lim_{N\to\infty}\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ exists and it is finite hence $\lim_{n\to\infty}W[h^{(n)}]-V[h^{*}]=0$.
+$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)-p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)\\ &\quad=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}\cdot h^{*}}{h^{(n)}\cdot Q_{Y|X}^{*}}\right)\\ &\quad\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{(n)}\log\left(\frac{h^{*}}{h^{(n)}}\right)-1=0.\end{split}$ (27)
+As a result, we obtain Since at any iteration $n$, $W[h^{(n)}]-V[h^{*}]\geq 0$ and for all integers $N>0$ the partial sum is upper-bounded by a constant $L(h^{*},h^{(0)})$ dependent only on the initial probability assignment $h^{(0)}$, we have that $\lim_{N\to\infty}\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ exists and it is finite hence $\lim_{n\to\infty}W[h^{(n)}]-V[h^{*}]=0$.
 This completes the proof.
@@
 We first derive the functional form of $J_{T}$:
-By defining matrices $M$, $\Gamma$ and $C$, respectively, as in (LABEL:matrix:_Newton_M), (14), (15), the matrix $J_{T}(\cdot)$ can be rewritten as in (13).
-To prove its invertibility, we need to ensure that 00 is not part of the set of eigenvalues of $J_{T}(y)$, i.e., $0\notin\operatorname{eig}(J_{T}(y)),~{}\forall y\in\mathbb{R}^{|\mathcal{X}|}$.
-Noticing that $M_{i,j}\geq 0,\forall(i,j)\in\mathcal{X}^{2}$ and $\sum_{i\in\mathcal{X}}M_{i,i}(y)=C_{i}(y)$, we can define the sets $D_{i}$ as:
-$D_{i}=\{\lambda\in\mathbb{R}:\left|\lambda-\left(C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\right)\right|\leq C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\}$ (28)
+By defining matrices $M$, $\Gamma$ and $C$, respectively, as in (14), (15), (16), the matrix $J_{T}(\cdot)$ can be rewritten as in (13).
+To prove its invertibility, we need to ensure that 0 is not part of the set of eigenvalues of $J_{T}(y)$, i.e., $0\notin\operatorname{eig}(J_{T}(y)),~\forall y\in\mathbb{R}^{|\mathcal{X}|}$.
+Noticing that $M[h^{(n)},u](i,j)\geq 0,\forall(i,j)\in\mathcal{X}^{2}$ and $\sum_{i\in\mathcal{X}}M[h^{(n)},u](i,j)=C[h^{(n)},u](i)$, we can define the sets $D_{i}$ as:
+$D_{i}=\Big{\{}\lambda\in\mathbb{R}:$ $\left|\lambda-\left(C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\right)\right|$
+$\leq C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\Big{\}}$
 Since $\bigcup_{i\in\mathcal{X}}D_{i}\subseteq\mathbb{R}_{0}^{+}$, we can apply Gershgorin Circle Theorem [47] to prove that $C[h^{(n)},u]-M[h^{(n)},u]$ has only non-negative eigenvalues.
@@
 Therefore, we obtain
-$J_{T}[h^{(n)},u]=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]\geq I>0$
+$J_{T}[h^{(n)},u]$ $=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]$
+$\geq I>0$
 This concludes the proof.
@@
 Let $V[\cdot]$, $W[\cdot]$ be the functionals defined in (24) and (25), respectively.
-Moreover, let $\hat{W}[\cdot]$ be a functional obtained by the alternating sequence $\hat{h}^{(n)}\to\hat{Q}_{Y|X}^{(n+1)}\to\hat{h}^{(n+1)}$ substituting $\hat{Q}_{Y|X}$ with fixed $\hat{h}$ as follows
-$\begin{split}\hat{W}[\hat{h}^{(n)}]=&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}\hat{h}^{(n)}(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}\end{split}$ (29)
-where $A^{(n)}=A[\hat{h}^{(n)}]$. Similarly to Theorem 2, we let $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and consider that
-$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{h}^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}\hat{h}^{(n)}(i)A^{(n)}(x,i,s)\right).\end{split}$
+Moreover, let $\hat{W}$$\cdot$] be a functional obtained by the alternating sequence $\hat{h}^{(n)}\to\hat{Q}_{Y|X}^{(n+1)}\to\hat{h}^{(n+1)}$ substituting $\hat{Q}_{Y|X}$ with fixed $\hat{h}$ as follows
+where $A^{(n)}=A[v^{(n)}]$. Similarly to Theorem 2, we let $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and consider that
+$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\\ &\quad=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{h}^{(n)}}\right)\\ &\qquad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\qquad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}\hat{h}^{(n)}(i)A^{(n)}(x,i,s)\right).\end{split}$
 Substituting the definition of $A(x,,s)$ and using (27), we obtain
-$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\leq\quad V[h^{*}]-\hat{W}[\hat{h}^{(n)}]\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}(p_{X}||h^{*})+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad\qquad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}.\end{split}$ (30)
+$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\leq\quad V[h^{*}]-\hat{W}[\hat{h}^{(n)}]\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}(p_{X}||h^{*})\\ &\qquad~~~~~~+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\qquad~~~~~~-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\qquad~~~~~~+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\\ &\qquad~~~~~~-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}.\end{split}$ (29)
 Note that the right-hand side of (29) can be bounded by
-$\begin{split}&\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=D_{f}\left(p_{X}||v^{(n)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{i\in\mathcal{X}}(h^{*}-v^{(n)})\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &\quad+D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||v^{(n)}\right)+\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &=-\left[D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),v^{(n)}}(h^{*})\right]+\left[D_{f}(p_{X}||h^{(n+1)})-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right]\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\end{split}$
+$\begin{split}&\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)\\ &\quad+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=D_{f}\left(p_{X}||v^{(n)}\right)-D_{f}\left(p_{X}||h^{*}\right)\\ &\quad+\sum_{i\in\mathcal{X}}(h^{*}-v^{(n)})\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &\quad+D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||v^{(n)}\right)\\ &\quad+\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=-\left[D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),v^{(n)}}(h^{*})\right]\\ &\quad+\left[D_{f}(p_{X}||h^{(n+1)})-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right]\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\end{split}$
 where $T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)$ is the first order Taylor expansion of $D_{f}(p_{X}||\cdot)$, centered in $v^{(n)}$ and evaluated in $h^{(n+1)}$, and $(a)$ is verified since $D_{f}(p_{X}||\cdot)$ is a convex function in its second argument.
@@
 where
-$\begin{split}G[\hat{h}^{(n)}]=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)+s_{P}\left[D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right].\end{split}$
+$\begin{split}&G[\hat{h}^{(n)}]=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)\\ &\qquad+s_{P}\left[D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right].\end{split}$
 Summing over $N$ terms we obtain
-$\begin{split}\sum_{n=1}^{N}\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)\\ &+s_{P}\sum_{n=1}^{N}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\\ \leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{\hat{Q}_{Y|X}^{(1)}}\right)+\sum_{n=1}^{N}o\left(||h^{(n+1)}-v^{(n)}||\right)\\ \leq&\quad\tilde{L}(h^{*},h^{(0)})\end{split}$
-where $\tilde{L}(h^{*},h^{(0)})$ is finite if the limit $\lim_{n\to\infty}||h^{(n+1)}-v^{(n)}||=0$ converges at least linearly. Thus, we can rewrite
+[…]
+[…]
 $0\leq\lim_{N\to\infty}\sum_{n=1}^{N}\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq\tilde{L}(h^{*},h^{(0)})$
@@
 $\mathcal{L}_{(D,P)}=\{Q_{Y|X}:\mathbb{E}_{Q_{Y|X}}\left[d(X,Y)\right]\leq D\land D_{f}(p_{X}||q_{Y})\leq P\}.$
-Furthermore, let $\Lambda_{s,v[q_{Y}]}$ be the set defined as
-$\Lambda_{s,v[q_{Y}]}\triangleq\big{\{}\lambda\in\mathbb{R}^{|\mathcal{X}|}:\forall x\in\mathcal{X},\lambda(x)\geq 0\land\forall y\in\mathcal{X},~{}\sum_{x\in\mathcal{X}}p_{X}(x)\lambda(x)A(x,,s)\leq 1\big{\}}$
+where $q_{Y}=\sum_{x\in\mathcal{X}}p_{X}Q_{Y|X}$. Furthermore, let $\Lambda_{s,v[q_{Y}]}$ be the set defined as
+$\Lambda_{s,v[q_{Y}]}\triangleq\Big{\{}\lambda\in\mathbb{R}^{|\mathcal{X}|}$ $\forall x\in\mathcal{X},\lambda(x)\geq 0\land\forall y\in\mathcal{X},$
+$\sum_{x\in\mathcal{X}}p_{X}(x)\lambda(x)A[v[q_{Y}]](x,y,s)\leq 1\Big{\}}.$
 Then, $\forall\mathbf{\lambda}\in\Lambda_{s,v[q_{Y}]}$, we obtain
-$\begin{split}R(D,P)\geq&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)-s_{D}D-s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y).\end{split}$
+$\begin{split}R(D,P)\geq&\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\lambda(x)\right)-s_{D}D\\ &-s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)g(p_{X}(y),v[q_{Y}](y)).\end{split}$
 Proof:
 Let $\mathbf{\lambda}\in\Lambda_{s,v[q_{Y}]}$ and $Q_{Y|X}\in\mathcal{L}_{(D,P)}$, then:
-$\begin{split}&I\left(p_{X},Q_{Y|X}\right)+\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)+s_{D}D+s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y)\\ &\geq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\log\left(\frac{Q_{Y|X}}{q_{Y}\lambda(x)A(x,,s)}\right)\\ &\geq 1-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\left(\frac{q_{Y}\lambda{x}A(x,,s)}{Q_{Y|X}}\right)\\ &=1-\sum_{y\in\mathcal{X}}q_{Y}\sum_{x\in\mathcal{X}}p_{X}\lambda(x)A(x,,s)\\ &\geq 1-\sum_{y\in\mathcal{X}}q_{Y}=0.\end{split}$
+$\begin{split}&I\left(p_{X},Q_{Y|X}\right)+\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)+s_{D}D\\ &\quad+s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)g(p_{X}(y),v[q_{Y}](y))\\ &\geq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\log\left(\frac{Q_{Y|X}}{q_{Y}\lambda(x)A[v[q_{Y}]](x,y,s)}\right)\\ &\geq 1-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\left(\frac{q_{Y}\lambda(x)A[v[q_{Y}]](x,y,s)}{Q_{Y|X}}\right)\\ &=1-\sum_{y\in\mathcal{X}}q_{Y}\sum_{x\in\mathcal{X}}p_{X}\lambda(x)A[v[q_{Y}]](x,y,s)\\ &\geq 1-\sum_{y\in\mathcal{X}}q_{Y}=0.\end{split}$
 The equality is ensured by the possibility of choosing $\lambda(x)\in\Lambda_{s,v[q_{Y}]}$ as
-$\begin{split}\lambda(x)=\frac{1}{\sum_{y\in\mathcal{X}}q_{Y}A(x,,s)}\\ \end{split}$
+[…]
 that once substituted describes the optimization problem found in Corollary 1. This completes the proof of the lemma. ∎
 Now we use Lemma 9 to prove Theorem 5. In particular, (19) can be derived from the following inequality:
-$\begin{split}R(D,P)&\leq\quad I\left(p_{X},\hat{Q}_{Y|X}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)\hat{Q}_{Y|X}(y|x)\log\left(\frac{\hat{Q}_{Y|X}(y,x)}{\sum_{x\in\mathcal{X}}p_{X}(x)\hat{Q}_{Y|X}(y|x)}\right)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+s_{P}\sum_{y\in\mathcal{X}}q_{Y}(y)\frac{p_{X}(y)}{v[q_{Y}](y)}\partial f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)\\ &\quad-s_{P}\Bigg{\{}\sum_{y\in\mathcal{X}}q_{Y}(y)f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)-P\Bigg{\}}-s_{P}\sum_{j\in\mathcal{X}}\hat{h}(j)A(x,j,s)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+\hat{W}[\hat{h}].\end{split}$
+[…]
 (18) is derived as an application of Lemma 9 by choosing $\lambda(x)$ as
-$\lambda(x)=\left(c_{\max}\sum_{y\in\mathcal{X}}\hat{h}(y)A(x,,s)\right)$
+$\lambda(x)=\left(c_{\max}\sum_{y\in\mathcal{X}}\hat{h}(y)A[v[\hat{h}]](x,y,s)\right)^{-1}$
 which respects the assumption of the theorem. This completes the proof.
@@
 The functional form of $J[h](i,j)$ in the case of Theorem 2 is obtained from
-$\begin{split}\frac{\partial S[h](i)}{\partial h(j)}&=h(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{\partial}{\partial h(j)}\left(\frac{A[S[h]](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s)}\right)+c[h,S[h]](i)\delta_{i,j}\end{split}$
+$\begin{split}&\frac{\partial S[h](i)}{\partial h(j)}=c[h,S[h]](i)\delta_{i,j}\\ &\quad+h(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{\partial}{\partial h(j)}\left(\frac{A[S[h]](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s)}\right)\end{split}$
 where
-$\begin{split}\frac{\partial(\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s))}{\partial h(j)}\ =\sum_{k\in\mathcal{X}}h(k)\frac{\partial A[S[h]](x,k,s)}{\partial h(j)}+A[S[h]](x,j,s)\end{split}$
-$\begin{split}\frac{\partial A(x,i,s)}{\partial h(j)}=-s_{P}A[S[h]](x,i,s)\frac{p_{X}(i)^{2}}{(S[h(i)])^{3}}\partial f\left(\frac{p_{X}(i)}{S[h(i)]}\right)\frac{\partial S[h](i)}{\partial h(j)}\end{split}.$
+$\begin{split}&\frac{\partial(\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s))}{\partial h(j)}=\\ &\qquad\sum_{k\in\mathcal{X}}h(k)\frac{\partial A[S[h]](x,k,s)}{\partial h(j)}+A[S[h]](x,j,s)\end{split}$
+$\begin{split}&\frac{\partial A(x,i,s)}{\partial h(j)}=\\ &\qquad-s_{P}A[S[h]](x,i,s)\frac{p_{X}(i)^{2}}{(S[h(i)])^{3}}\partial f\left(\frac{p_{X}(i)}{S[h(i)]}\right)\frac{\partial S[h](i)}{\partial h(j)}\end{split}.$
 By noticing that
 $\frac{p_{X}(i)^{2}}{(S[h](i))^{3}}\partial f\left(\frac{p_{X}(k)}{S[h](i)}\right)=\frac{\partial^{2}}{\partial q(i)^{2}}D_{f}(p_{X}||q)\Big{|}_{S[h]}$
-and defining the matrices $M$ and $\Gamma$ as in (LABEL:matrix:_Newton_M) and (14), we can rewrite the entries of $J[h^{*}]$ as
-$\begin{split}J[h^{*}](i,j)&=c[h^{*},h^{*}](i)\left(\delta_{i,j}-\Gamma[h^{*},h^{*}](i)J[h^{*}](i,j)\right)\\ &\quad+\sum_{k\in\mathcal{X}}\Gamma[h^{*},h^{*}](k)M[h^{*},h^{*}](i,k)J[h^{*}](k,j)-M[h^{*},h^{*}](i,j)\\ \end{split}$ (31)
+and defining the matrices $M$ and $\Gamma$ as in (14) and (15), we can rewrite the entries of $J[h^{*}]$ as
+$J[h^{*}](i,j)$ $=c[h^{*},h^{*}](i)\left(\delta_{i,j}-\Gamma[h^{*},h^{*}](i)J[h^{*}](i,j)\right)$
+$\quad+\sum_{k\in\mathcal{X}}\Gamma[h^{*},h^{*}](k)M[h^{*},h^{*}](i,k)J[h^{*}](k,j)$
+$\quad-M[h^{*},h^{*}](i,j)$ (30)
 where $\delta_{i,j}$ is the Kronecker delta. As a result, we can express (30) in matrix form as follows
-$\begin{split}J[h^{*}]&=C[h^{*},h^{*}](I-\Gamma[h^{*},h^{*}]J[h^{*}])-M[h^{*},h^{*}]+M[h^{*},h^{*}]\Gamma[h^{*},h^{*}]J[h^{*}]\\ &=(C[h^{*},h^{*}]-M[h^{*},h^{*}])(I-\Gamma J[h^{*}])\end{split}$
+$\begin{split}J[h^{*}]&=C[h^{*},h^{*}](I-\Gamma[h^{*},h^{*}]J[h^{*}])\\ &\quad-M[h^{*},h^{*}]+M[h^{*},h^{*}]\Gamma[h^{*},h^{*}]J[h^{*}]\\ &=(C[h^{*},h^{*}]-M[h^{*},h^{*}])(I-\Gamma J[h^{*}])\end{split}$
 where $C[\cdot,\cdot]$ is defined in (16). Finally, we obtain (21) noticing that $C[h^{*},h^{*}]=I$ due to the optimality conditions found in Lemma 2, thus concluding the proof.
@@
 where
-$\begin{split}\sum_{i\in\mathcal{X}}M^{*}(i,j)&=\sum_{i\in\mathcal{X}}h^{*}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,i,s)A[h^{*}](x,j,s){}}{(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s))^{2}}\\ &=\sum_{x\in\mathcal{X}}\frac{A[h^{*}](x,j,s)}{\left(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s)\right)^{2}}\left(\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)\right)\\ &=c[h^{*},h^{*}](j)=1.\end{split}$
+$\sum_{i\in\mathcal{X}}M^{*}(i,j)$
+$=\sum_{i\in\mathcal{X}}h^{*}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,i,s)A[h^{*}](x,j,s){}}{(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s))^{2}}$
+$=\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,j,s)\left(\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)\right)}{\left(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s)\right)^{2}}$
+$=c[h^{*},h^{*}](j)=1.$
 Thus the eigenvalues of $M^{*}$ are each in at least one of the disks $I_{i}=\{z\in\mathbb{R}:|z-M^{*}(i,i)|\leq 1-M^{*}(i,i)\}$, which are all contained in the disk $I=\{z\in\mathbb{R}:|z|\leq 1\}$.
@@
 Using Lemmas 5 and 6, the following inequalities hold
-$1\displaystyle 0<\operatorname{eig}(M^{*})\leq 1\implies 0\leq\operatorname{eig}(I-M^{*})<$ 1
+$0<\operatorname{eig}(M^{*})\leq 1$ $\implies 0\leq\operatorname{eig}(I-M^{*})<1$
 $0\leq\operatorname{eig}(I-M^{*})<1$ $\xRightarrow{(a)}1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)$
-$1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)\implies 0<\operatorname{eig}((I+(I-M^{*})\Gamma^{*})^{-1})\leq 1$
+$1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)$ $\implies$
+$0<\operatorname{eig}((I+(I-M^{*})\Gamma^{*})^{-1})\leq 1$
 where $(a)$ is due to $\Gamma$ being a positive definite matrix. Using the previous inequalities, we can rewrite (21) as follows:
@@
 Define $\theta_{\sup}\triangleq\operatorname{eig}_{\max}(I-M^{*})\cdot\operatorname{eig}_{\max}((I+(I-M^{*})\Gamma^{*})^{-1})$.
-Then, we can show that $10\leq\operatorname{eig}(J[h^{*}])\leq\theta_{\sup}<$ 1 is always verified.
-The second part of the theorem follows directly from [42, Theorem 5] hence we omit it.
+Then, we can show that $0\leq\operatorname{eig}(J[h^{*}])\leq\theta_{\sup}<1$ is always verified.
+The second part of the theorem follows directly from [44, Theorem 5] hence we omit it.
 This completes the proof.
@@
 $\begin{split}(I-\Gamma J(h^{*}))\geq 0\implies s_{P,\max}\leq\frac{1}{\theta_{\max}f^{\prime\prime}(0)}.\end{split}$
-In order to guarantee $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1, it is sufficient to have
+In order to guarantee $0\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<1$, it is sufficient to have
 $\begin{split}(I-\Gamma)\geq 0\implies s_{P,\max}\leq\frac{1}{f^{\prime\prime}(0)}.\end{split}$
-Since in the non-degenerate case we have $10<\theta_{\max}<$ 1, we can construct $\epsilon^{\prime}=\min\left\{s_{P,\max},\frac{1}{f^{\prime\prime}(0)}\left(\frac{1}{\theta_{\max}}-1\right)\right\}$ > 0 and define the set $I_{s_{P}}^{\epsilon^{\prime}}$ where the exponential convergence of the approximate algorithm is guaranteed.
+Since in the non-degenerate case we have $0<\theta_{\max}<1$, we can construct $\epsilon^{\prime}=\min\left\{s_{P,\max},\frac{1}{f^{\prime\prime}(0)}\left(\frac{1}{\theta_{\max}}-1\right)\right\}$ > 0 and define the set $I_{s_{P}}^{\epsilon^{\prime}}$ where the exponential convergence of the approximate algorithm is guaranteed.
 This concludes the proof.
 Appendix L Proof of Lemma 8
-We remind that $f=|x-1|$ is the function associated with the TV distance $TV(\cdot||\cdot)$.
+We remind that f=|x1|f=|x-1| is the function associated with the TV distance $TV(\cdot||\cdot)$.
 The first statement can be proved by first establishing the uniform convergence of $f_{n}\to f$ as $n\to\infty$
-$\sup_{x\in\mathbb{R}}\left|f_{n}(x)-f(x)\right|$ $=\sup_{x\in\mathbb{R}}\left|\frac{2}{\pi}(x-1)\arctan(n(x-1))-|x-1|\right|$
+$\sup_{x\in\mathbb{R}}\left|f_{n}(x)-f(x)\right|$
+$=\sup_{x\in\mathbb{R}}\left|\frac{2}{\pi}(x-1)\arctan(n(x-1))-|x-1|\right|$
 $=\sup_{x\in\mathbb{R}^{+}_{0}}\left|\frac{2}{\pi}x\arctan(nx)-x\right|$
 $=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\left(\frac{\pi}{2}-\arctan(nx)\right)$
-$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\arctan\left(\frac{1}{nx}\right)$
-$\leq\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{n\pi}=\frac{2}{n\pi}$
+$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\arctan\left(\frac{1}{nx}\right)\leq\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{n\pi}=\frac{2}{n\pi}$
 meaning that $\lim_{n\to\infty}\sup_{x\in\mathbb{R}}|f_{n}(x)-f(x)|=0$.
 A direct consequence of the above is that $D_{f_{n}}\to TV$ uniformly in the limit of $n\to\infty$ since, for any $p,q\in\mathcal{P}(\mathcal{X})$,
-$\lim_{n\to\infty}|D_{f_{n}}(p||q)-TV(p||q)|$ $=\lim_{n\to\infty}\left|\sum_{x\in\mathcal{X}}q(x)\left(f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right)\right|$
+$\lim_{n\to\infty}|D_{f_{n}}(p||q)-TV(p||q)|$
+$=\lim_{n\to\infty}\left|\sum_{x\in\mathcal{X}}q(x)\left(f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right)\right|$
 $\leq\sum_{x\in\mathcal{X}}q(x)\lim_{n\to\infty}\left|f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right|=0.$